### v2

# LLMs may Dominate Information Access: Neural Retrievers are Biased Towards LLM-Generated Texts

Abstract.
Recently, the emergence of large language models (LLMs) has revolutionized the paradigm of information retrieval (IR) applications, especially in web search. With their remarkable capabilities in generating human-like texts, LLMs have created enormous texts on the Internet. As a result, IR systems in the LLMs era are facing a new challenge: the indexed documents now are not only written by human beings but also automatically generated by the LLMs. How these LLM-generated documents influence the IR systems is a pressing and still unexplored question. In this work, we conduct a quantitative evaluation of different IR models in scenarios where both human-written and LLM-generated texts are involved. Surprisingly, our findings indicate that neural retrieval models tend to rank LLM-generated documents higher. We refer to this category of biases in neural retrieval models towards the LLM-generated text as the source bias. Moreover, we discover that this bias is not confined to the first-stage neural retrievers, but extends to the second-stage neural re-rankers. Then, we provide an in-depth analysis from the perspective of text compression and observe that neural models can better understand the semantic information of LLM-generated text, which is further substantiated by our theoretical analysis. To mitigate the source bias, we also propose a plug-and-play debiased constraint for the optimization objective, and experimental results show the effectiveness. Finally, we discuss the potential severe concerns stemming from the observed source bias and hope our findings can serve as a critical wake-up call to the IR community and beyond. To facilitate future explorations of IR in the LLM era, the constructed two new benchmarks and codes will later be available at https://github.com/KID-22/LLM4IR-Bias.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: XXXX; August 03–05, 2023; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
(a) IR in the Pre-LLM Era
(b) IR in the LLM Era
Figure 1. The overview evolution of IR paradigm from the Pre-LLM era to the LLM era.
With the advent of large language models (LLMs), exemplified by ChatGPT, the field of artificial intelligence generated content (AIGC) has surged to new heights of prosperity (Cao et al., 2023; Wu et al., 2023a). LLMs have demonstrated their remarkable capabilities in automatically generating human-like text at scale, resulting in the Internet being inundated with an unprecedented volume of AIGC content (Wei et al., 2022; Spitale et al., 2023). This influx of LLM-generated content has fundamentally reshaped the digital ecosystem, challenging conventional paradigms of content creation, dissemination, and information access on the Internet (Ai et al., 2023; Zhu et al., 2023).
Meanwhile, information retrieval (IR) systems have become the keystone of how individuals navigate and access information within the vast expanse of the Internet (Singhal et al., 2001; Manning, 2009). The core function of an IR system is to efficiently retrieve relevant documents from the target corpus in response to users’ queries (Liu et al., 2009; Li, 2022; Xu and Li, 2007). As illustrated in Figure 1, in the era preceding the widespread emergence of LLMs, IR systems primarily revolved around the retrieval of relevant documents solely from the human-written corpus. However, with the rise of LLMs, IR systems are confronted with a profound paradigm shift. That is, as LLMs have significantly driven the proliferation of AIGC, the corpus of IR systems now gradually includes both human-written and LLM-generated texts. Consequently, this paradigm shift raises a fundamental research question: What is the impact of the proliferation of generated content on IR systems? In this LLM era, we aim to explore whether existing retrieval models tend to prioritize LLM-generated text over human-written text, even when both convey similar semantic information. If this holds, LLMs may dominate information access, particularly as their generated content is rapidly growing on the Internet (Hanley and Durumeric, 2023).
To approach the fundamental research question, we decompose it into three specific research questions. The first question is RQ1: How to construct a simulation environment to evaluate IR models in the LLM era? As the LLM-generated content is currently unidentifiable on the Internet, there are no public retrieval benchmarks where the corpus contains both human-written text and LLM-generated text. To construct a more realistic evaluation environment, we introduce a natural approach to synthesize LLM-generated data without incurring significant costs for human annotation of relevant labels. Specifically, we leverage the original human-written texts as the instruction conditions to prompt LLMs to generate rewritten text copies while preserving the same semantic meaning. In this way, we can confidently assign the relevant labels for LLM-generated data. Extensive empirical analysis validates the quality of our constructed environment, as it can accurately reflect real-world situations of IR in the LLM era. As a result, we introduce two new benchmarks, SciFact+AIGC and NQ320K+AIGC, tailored for IR research in the LLM era.
Based on the constructed environment, we can explore further RQ2: Are and where are retrieval models biased towards LLM-generated texts? We conduct systematical experiments with various representative and state-of-the-art retrieval models, ranging from traditional lexical models to modern neural retrieval models (Zhao et al., 2022; Yates et al., 2021; Guo et al., 2022). Surprisingly, our findings indicate that neural retrieval models are biased towards LLM-generated texts, i.e., rank LLM-generated texts in a higher position. We refer to this as source bias, as the neural retrievers favor texts from specific sources (i.e., LLM-generated). Our further experiments indicate that the source bias not only extends to the second-stage neural re-rankers but also manifests in a more severe manner. These findings further corroborate the prevalence of source bias in neural ranking models that rely on pretrained language models (PLMs) (Yates et al., 2021; Guo et al., 2020; Zhao et al., 2022). In other words, these neural ranking models appear to adopt some inherent semantic shortcuts to learn the semantic representation of documents generated by LLM.
Then, what we are curious about is RQ3: Why are neural retrieval models biased towards LLM-generated texts? Inspired by the recent studies that posit LLMs are lossless compressors (Delétang et al., 2023), we aim to analyze the cause of source bias from the viewpoint of text compression. We first examine the singular value (Klema and Laub, 1980) of different corpora and uncover that LLM-generated text typically maintains more focused semantics with minimal noise, rendering it better understood for precise semantic matching. Furthermore, to better examine whether the neural retrieval model based on PLMs can better understand the LLM-generated texts, we conduct an in-depth analysis for perplexity. Our experimental and theoretical analysis suggest that LLM-generated text consistently yields lower perplexity scores, signifying greater comprehensibility and higher confidence from the perspective of PLMs. These observations confirm that LLM-generated text is more readily understandable to neural retrieval models, thereby resulting in source bias.
As our analysis has revealed source bias, we aim to further study the RQ4: How can we mitigate the source bias in neural retrieval models? Specifically, we propose a simple but effective debiased constraint to penalize biased samples during optimization. In this way, we can force the retrieval models to focus on modeling semantic relevance rather than the inherent shortcut between neural retrievers and LLM-generated content, leading to the mitigation of source bias. This debiased constraint can be plugged and played into the ranking optimization objective of various neural retrieval models. Besides, our approach offers the capability to control the degree of bias removal, providing flexibility in determining whether the two sources of content should be treated equally or biased toward human-written content according to specific requirements and environmental considerations.
Finally, we discuss the potential emerging concerns stemming from source bias. First, the presence of source bias poses a significant risk of gradually rendering human-written content less accessible, potentially causing a disruption in the content ecosystem. This concern is further compounded by the growing prevalence of LLM-generated text on the Internet(Hanley and Durumeric, 2023; Bengio et al., 2023). Second, there is the risk that source bias may amplify the spread of misinformation, especially considering the potential of LLMs to generate deceptive content, whether intentionally or not (Chen and Shu, 2023; Su et al., 2023). Third, source bias may be maliciously exploited to manipulate existing retrieval and ranking algorithms, creating a precarious vulnerability that could be weaponized by malicious actors. In light of these pressing issues, we hope that our findings serve as a resounding wake-up call to all stakeholders involved in IR systems and beyond.
In summary, the contributions of this paper are as follows:
(1) This is the first work that introduces the paradigm of IR systems in the LLM era, where the retrieval corpus consists of both human-written texts and LLM-generated texts. This scenario is a more realistic setting, considering the growing prosperity of AIGC.
(2) We uncover the source bias in the LLM era, a new inherent bias in neural retrieval models, where both the neural retrievers and re-rankers rank the LLM-generated text in a higher position than human-written text with the almost same semantic relevance.
(3) We provide an in-depth analysis and theoretical insights of source bias from the viewpoint of text compression. The analysis indicates that LLM-generated text maintains more focused semantics with minimal noise and is more readily comprehensible for neural retrieval models.
(4) We propose a debiased constraint to penalize the bias of neural retrieval models during the optimization and experimental results demonstrate the effectiveness in mitigating source bias.
(5) We also provide two new benchmarks, SciFact+AIGC and NQ320K+AIGC, which contain both high-quality human-written and various LLM-generated corpus and corresponding relevant labels. We believe these two benchmarks can serve as valuable resources for facilitating future research of IR in the LLM era.
Figure 2. The overall paradigm of the proposed evaluation framework for IR in the LLM era.
2. RQ1: Environment Construction
With the increasing usage of LLMs in generating texts (e.g., paraphrasing or rewriting texts with LLMs), the corpus of IR systems includes both human-written and LLM-generated texts nowadays. In this section, we introduce a novel and natural framework for the quantitative evaluation of retrieval models in the LLM era, i.e., in scenarios where the corpus includes both human-written and LLM-generated texts, as shown in Figure 2. To better align with the real-world scenarios in the LLM era, the simulated environments should meet the following two essential criteria. Firstly, it is imperative to distinguish between human-written texts and LLM-generated texts within the corpus. Secondly, we need access to pertinent labels for LLM-generated data in response to queries. To ensure the most effective evaluation of retrieval model preferences, every piece of human-written text should have a corresponding LLM-generated counterpart with consistent content.
2.1. Preliminary
2.1.1. Notation
Formally, in the Pre-LLM era, given a query text $q\in\mathcal{Q}$ where $\mathcal{Q}$ is the set of all queries, the traditional IR system aims to retrieve a list of top-$K$ relevant documents $\{d^{(1)},d^{(2)},\ldots,d^{(K)}\}$ from a corpus $\mathcal{C}^{H}=\{d_{1}^{H},d_{2}^{H},\ldots d_{N}^{H}\}$ which consists of $N$ human-written documents. However, in the era of LLMs, there is also LLM-generated text in the corpus. To simulate and evaluate the IR models in the LLM era, we also create an additional corpus $\mathcal{C}^{G}=\{d_{1}^{G},d_{2}^{G},\ldots,d_{N}^{G}\}$ where each document is generated by a LLM, e.g., $d_{1}^{G}$ can be created by ChatGPT by constructing a prompt that asks ChatGPT to rewrite $d_{1}^{H}$ while preserving its original meaning. Consequently, given a user query $q$, the primary objective of a retriever in the LLM era is to return the top-$K$ relevant documents from the combined corpus $\mathcal{C}=\mathcal{C}^{H}\bigcup\mathcal{C}^{G}$.
2.1.2. Retrieval Models
IR models have witnessed a significant evolution, shifting from traditional lexical approaches primarily reliant on term matching to the recent neural IR models founded on semantic matching (Zhao et al., 2022; Guo et al., 2022; Yates et al., 2021; Gonzalo et al., 2014). In this study, we select the following representative retrieval models to conduct a thorough exploration of their performance in scenarios encompassing both human-written and LLM-generated texts:
Lexical Retrieval Models are based on the exact term matching between high-dimensional sparse vectors of query and document. We choose the widely used two approaches TF-IDF (Sparck Jones, 1972) and BM25 (Robertson et al., 2009) in our experiments.
Neural Retrieval Models utilize neural networks to learn semantic relationships between query and document to improve the lexical gap (Li et al., 2014; Zhao et al., 2022; Yates et al., 2021; Guo et al., 2020). We select the following four representative neural retrieval models: ANCE (Xiong et al., 2020) is a dense retrieval model that selects hard training negatives globally from the entire corpus, using an asynchronously updated ANN index. BERM (Xu et al., 2023) is also a dense retrieval model that focuses on learning the domain-invariant features in information retrieval tasks. TAS-B (Hofstätter et al., 2021) is a bi-encoder trained with a balanced margin sampling technique that samples queries out of a cluster per batch, ensuring efficiency in the sampling process. Contriever (Izacard et al., 2021) uses momentum contrastive learning to train dense retrievers from text paragraphs and obtain state-of-the-art zero-shot retrieval performance.
Table 1. Statistics of the constructed two datasets. Avg. Doc / Query means the average number of relevant documents per query.
Dataset # Test Queries # Avg. Query Length Human-Written Corpus Llama2-Generated/ChatGPT-Generated Corpus
# Corpus Avg. Doc Length Avg. Doc / Query # Corpus Avg. Doc Length Avg. Doc / Query
SciFact+AIGC 300 12.38 5,183 201.81 1.1 5,183/5,183 192.66/203.57 1.1/1.1
NQ320K+AIGC 7,830 9.24 109,739 199.79 1.0 109,739/- 174.49/- 1.0/-
2.2. Constructing IR Datasets in the LLM Era
To better align with real-world IR scenarios in the LLM era, we first prompt LLMs to rewrite human-written corpus to build two new standard retrieval datasets: SciFact+AIGC and NQ320K+AIGC. These two new datasets can serve as valuable resources to facilitate future research of IR in the LLM era.
(a) Jaccard Similarity
(b) Overlap
Figure 3. Distribution of term Jaccard similarity and overlap between Llama2-generated and human-written corpora.
(a) SciFact+AIGC
(b) NQ320K+AIGC
Figure 4. Semantic embedding visualization of different corpora on SciFact+AIGC and NQ320K+AIGC datasets.
2.2.1. Human-Written Corpus
We first choose two widely used retrieval datasets written by humans in the Pre-LLM era as the seed data: SciFact and NQ320K. SciFact [^1] https://allenai.org/data/scifact (Wadden et al., 2020) dataset aims to retrieve evidence from the research literature containing scientific paper abstracts for fact-checking. NQ320K [^2] https://ai.google.com/research/NaturalQuestions (Kwiatkowski et al., 2019) is based on the Natural Questions (NQ) dataset from Google, where the documents are gathered from Wikipedia pages, and the queries are natural language questions. Following the practice in BEIR benchmark (Thakur et al., 2021), we process these two datasets in a standard format: corpus $\mathcal{C}^{H}$, queries $\mathcal{Q}$, and labels $\mathcal{R}^{H}=\{(q_{m},d^{H}_{m},r_{m})\}^{M}_{m=1}$, where $M$ is the number of labeled query-document pairs in the corpus.
2.2.2. LLM-Generated Corpus
Constructing an IR dataset in the LLM era typically involves two steps: collecting the LLM-generated corpus $\mathcal{C}^{G}$ and employing human evaluators from crowdsourcing platforms to annotate relevance labels for each query-document pair. Given the significant cost associated with human annotation, we adopt an alternative natural approach. Instead, we repurpose the original human-written corpus as our seed data source and directly instruct LLMs to rewrite each given text from the human-written corpus. As the written text generated by LLM carries almost the same semantic information as the original human-written text, we can assign the same relevance labels to new query-document pairs as those assigned to the original query-document pairs.
Our instruction is straightforward: “Please rewrite the following text: {{human-written text}}”, as illustrated in the left part of Figure 2. This straightforward instruction enables LLMs to generate text without too many constraints while maintaining semantic equivalence to the original human-written text. Specifically, we choose Llama2 (Touvron et al., 2023) and ChatGPT to rewrite each seed human-written corpus, as Llama2 and ChatGPT are both the most widely-used and nearly the state-of-the-art open-sourced and closed-source LLM, respectively. We only generate texts with ChatGPT corresponding to the texts in SciFact dataset, mainly due to the significant cost involved in processing the larger NQ320K dataset. If not specified, the temperature of the used LLM for generation corpus is set as 0.2 by default for the following experiments.
As a result, we can obtain two corresponding LLM-generated corpora with SciFact and NQ320K as seed data. After that, we extend the original labels of query and human-written text $\mathcal{R}^{H}=\{(q_{m},d^{H}_{m},r_{m})\}^{M}_{m=1}$ to get the corresponding label of LLM-generated text $\mathcal{R}^{G}=\{(q_{m},d^{G}_{m},r_{m})\}^{M}_{m=1}$. We will validate the quality of the datasets in the following section. Combining each original human-written corpus $\mathcal{C}^{H}$ with its corresponding LLM-generated corpus $\mathcal{C}^{G}$, original queries $\mathcal{Q}$, and labels $\mathcal{R}^{H}\bigcup\mathcal{R}^{G}$, we can create two new datasets, denoted as SciFact+AIGC and NQ320K+AIGC.  Table 1 summarize the statistics of the proposed two datasets.
[…]
Model Model Corpus SciFact+AIGC NQ320K+AIGC
Type NDCG@1 NDCG@3 NDCG@5 MAP@1 MAP@3 MAP@5 NDCG@1 NDCG@3 NDCG@5 MAP@1 MAP@3 MAP@5
Lexical TF-IDF Human-Written 42.0 49.5 52.7 40.7 47.1 49.0 12.2 15.8 16.8 12.2 14.9 15.5
LLM-Generated 43.0 49.8 52.6 40.8 47.5 49.2 9.4 12.6 13.9 9.4 11.8 12.5
BM25 Human-Written 46.0 54.2 56.3 43.8 51.5 52.8 12.9 16.3 17.6 12.9 15.5 16.2
LLM-Generated 46.3 53.6 55.3 44.1 51.1 52.2 11.9 15.3 16.5 11.9 14.5 15.1
Neural ANCE Human-Written 38.7 44.3 46.5 36.3 41.9 43.3 50.6 60.0 62.2 50.6 57.7 58.9
LLM-Generated 41.0 46.0 48.2 37.8 43.5 45.0 49.3 58.8 61.2 49.3 56.5 57.8
BERM Human-Written 37.0 42.1 44.2 34.7 39.7 41.3 49.2 58.3 60.4 49.2 56.1 57.3
LLM-Generated 40.7 44.5 46.2 37.7 42.3 43.5 48.4 57.5 59.8 48.4 55.3 56.5
TAS-B Human-Written 52.7 58.1 60.2 49.9 55.6 57.2 53.4 63.0 65.4 53.4 60.7 62.0
LLM-Generated 50.7 57.0 58.9 48.0 54.6 55.9 51.9 62.3 64.7 51.9 59.8 61.1
Contriever Human-Written 54.0 61.8 63.2 51.4 58.9 60.0 58.2 68.4 70.3 58.2 65.9 67.0
LLM-Generated 55.7 62.0 64.8 52.9 59.5 61.5 57.1 67.5 69.8 57.1 64.9 66.2
2.3. Statistics and Quality Validation of Datasets
Take the Llama2-generated data as an example, we conduct the statistics and quality validation of the constructed datasets. The analysis of ChatGPT-generated datasets is provided in Appendix A.4, which shows similar observations and conclusions.
2.3.1. Term-based Statistics and Analysis
We first analyze the term-based similarity between the LLM-generated corpus and the human-written corpus. Specifically, we compute the Jaccard similarity ($\frac{|d^{G}\bigcap d^{H}|}{|d^{G}\bigcup d^{H}|}$) and the overlap ($\frac{|d^{G}\bigcap d^{H}|}{|d^{H}|}$) between each LLM-generated document and orginal human-written document. The distribution of these two term-based similarity metrics is shown in Figure 3. As we can see, both the Jaccard similarity and overlap distributions exhibit normal distribution, with peaks at about 0.4 and 0.6 for NQ320K+AIGC, and about 0.6 and 0.8 for SciFact+AIGC, respectively. These observations suggest that while there is a considerable overlap of terms between the LLM-generated text and the original human-written text, there are also distinct differences, especially noticeable in the NQ320K+AIGC dataset.
Figure 5. Distribution of cosine similarity of semantic embedding between Llama2-generated and human-written corpora.
2.3.2. Semantic-based Statistics and Analysis
For the LLM-generated texts, a pivotal consideration is whether they faithfully preserve the underlying semantics of the corresponding human-written corpus. If they indeed do so, we can confidently assign them the same relevance labels as the labels of their corresponding original human-written texts given each query.
To assess this, we first leverage the OpenAI embedding model333text-embedding-ada-002:https://platform.openai.com/docs/api-reference/embeddings to acquire semantic embeddings for both the LLM-generated and human-written corpora. Subsequently, we visualize these embeddings through T-SNE (Van der Maaten and Hinton, 2008) in Figure 4. We observe a strikingly close overlap between the Llama2-generated corpus with the human-written corpus in the latent space. This observation strongly suggests that these Llama2-generated corpora adeptly preserve the original semantics. Moreover, we delve into the cosine similarity of semantic embeddings between the LLM-generated text and their corresponding human-written counterparts. The results, as shown in Figure 5, also indicate a high degree of similarity, with most values exceeding 0.95, affirming the faithful preservation of semantics in LLM-generated text. Hence, for each query-document pair $(q,d^{G})$, we can confidently assign the relevant label $r$ to be the same as that of $(q,d^{H})$.
To further validate the accuracy of the relevance label assignments, we conduct an evaluation of retrieval models on the human-written corpus and the LLM-generated corpus, respectively. The results are presented in Table 2. It is evident that all retrieval models exhibit no significant performance discrepancies in terms of various ranking metrics between the human-written and LLM-generated corpora across all datasets. This observation reinforces the confidence in the quality of our newly constructed datasets.
Table 3. Performance comparison of retrieval models for mixed human-written and Llama2-generated corpora on SciFact+AIGC and NQ320K+AIGC dataset. The numbers indicate that retrieval models rank human-written documents in higher positions than LLM-generated documents (i.e., $\text{Relative}~{}~{}\Delta>0\%$). Conversely, the numbers mean retrieval models rank LLM-generated documents in higher positions than human-written documents (i.e., $\text{Relative}~{}~{}\Delta\leq 0\%$). The intensity of the color reflects the extent of the difference. In the subsequent tables, we will continue with this color scheme.
Model Model Target Corpus SciFact+AIGC NQ320K+AIGC
Type NDCG@1 NDCG@3 NDCG@5 MAP@1 MAP@3 MAP@5 NDCG@1 NDCG@3 NDCG@5 MAP@1 MAP@3 MAP@5
Lexical TF-IDF Human-Written 22.0 36.9 39.7 21.2 33.0 34.7 7.1 11.0 12.3 7.1 10.0 10.8
LLM-Generated 17.0 33.8 37.2 16.2 29.5 31.5 3.4 8.1 9.4 3.4 7.0 7.7
$\text{Relative}~{}~{}\Delta$ 25.6 8.8 6.5 26.7 11.2 9.7 70.5 30.4 26.7 70.5 35.3 33.5
BM25 Human-Written 26.7 40.3 44.4 25.7 36.7 39.1 7.2 11.6 12.9 7.2 10.6 11.3
LLM-Generated 21.0 38.8 41.5 19.6 34.3 35.9 6.1 10.9 11.9 6.1 9.7 10.3
$\text{Relative}~{}~{}\Delta$ 23.9 3.8 6.8 26.9 6.8 8.5 16.5 6.2 8.1 16.5 8.9 9.3
Neural ANCE Human-Written 15.3 30.1 32.7 14.2 26.2 27.7 22.2 41.2 44.6 22.2 36.9 38.8
LLM-Generated 24.7 35.8 37.7 23.3 32.4 33.6 29.1 45.9 49.0 29.1 42.0 43.8
$\text{Relative}~{}~{}\Delta$47.0 -17.3 -14.2 -48.5 -21.2 -19.2 -26.9 -10.8 -9.4 -26.9 -12.9 -12.1
BERM Human-Written 16.3 30.2 31.8 15.7 26.5 27.5 18.6 37.5 40.7 18.6 33.1 34.9
LLM-Generated 23.7 34.1 36.4 21.7 30.8 32.2 31.6 47.0 50.0 31.6 43.5 45.1
$\text{Relative}~{}~{}\Delta$37.0 -12.1 -13.5 -32.1 -15.0 -15.7 -51.8 -22.5 -20.5 -51.8 -27.2 -25.5
TAS-B Human-Written 20.0 40.2 43.1 19.5 35.2 36.9 25.7 45.4 48.8 25.7 40.9 42.8
LLM-Generated 31.7 44.8 47.5 29.7 41.1 42.7 27.6 46.5 50.0 27.6 42.2 44.2
$\text{Relative}~{}~{}\Delta$45.3 -10.8 -9.7 -41.5 -15.5 -14.6 -7.1 -2.4 -2.4 -7.1 -3.1 -3.2
Contriever Human-Written 24.0 43.7 47.8 23.3 38.8 41.2 25.9 48.5 51.9 25.9 43.3 45.3
LLM-Generated 31.0 47.8 50.5 29.6 43.2 44.8 32.5 51.9 55.4 32.5 47.5 49.4
$\text{Relative}~{}~{}\Delta$25.5 -9.0 -5.5 -23.8 -10.7 -8.4 -22.6 -6.8 -6.5 -22.6 -9.3 -8.7
3. RQ2: Explore The Source Bias
In this section, we conduct extensive experiments on the constructed datasets to explore the source bias from various aspects. With the constructed simulated environment, we first introduce the evaluation metrics to quantify the severity of source bias. We then conduct experiments with different retrieval models on both the first-stage retrieval and the second-stage re-ranking. Finally, we also conduct experiments on corpus generated by LLM with different temperatures to further investigate the source bias.
3.1. Evaluation Metrics for Source Bias
To quantitatively explore source bias, we calculate the ranking metrics, targeting separately either human-written or LLM-generated corpus. Specifically, for each query, an IR model produces a ranking list that comprises documents from both human-written and LLM-generated corpora. We then calculate ranking metrics separately for human-written and LLM-generated texts, depending on the target data source. When computing metrics for one target (e.g., human-written corpus $\mathcal{C}^{H}$), the data corresponding to the other side (e.g., LLM-generated corpus $\mathcal{C}^{G}$) is masked or disregarded (i.e., mask all the positive label $r\in\mathcal{R}^{G}$ as negative), but the rank of each document is still based on the original ranking list that incorporates a mixture of both types of text. This approach allows us to independently assess the performance of IR models on different corpus sources. For the ranking metrics, we report the top-$K$ Normalized Discounted Cumulative Gain (NDCG@$K$) and Mean Average Precision (MAP@$K$) with $K\in\{1,3,5\}$. For a better understanding of the calculation of the evaluation metrics, we provide a toy example in Appendix A.1.
To better normalize the difference among different benchmarks, we also introduce the relative percentage difference as follows:
Relative $\text{Relative $\Delta$}=\frac{\texttt{Metric}_{\text{Human-written}}-\texttt{Metric}_{\text{LLM-generated}}}{\frac{1}{2}(\texttt{Metric}_{\text{Human-written}}+\texttt{Metric}_{\text{LLM-generated}})}\times 100\%,$
where the Metric can be NDCG@$K$ and MAP@$K$. Note that Relative $\Delta>0$ means retrieval models rank human-written texts higher, and Relative $\Delta<0$ indicates LLM-generated texts are ranked higher. The greater the absolute value of Relative $\Delta$, the greater the ranking performance difference on two corpora.
3.2. Bias in Different Retrieval Models
Our initial set of experiments is aimed at conducting a comprehensive analysis of different retrieval models within the context of the LLM era. These models encompass both traditional lexical approaches and modern neural IR models. The results targeted on either human-written or LLM-generated corpus on the SciFact+AIGC and NQ320K+AIGC are reported in Table 3. Based on these results, we have the following noteworthy observations and conclusions:
Lexical approaches tend to favor human-written documents. When the retrieved corpora contain both human-written and LLM-generated texts, TF-IDF and BM25 achieve better results on the human-written corpus. This phenomenon is observed in terms of almost all ranking metrics across both the SciFact+AIGC and NQ320K+AIGC datasets. A plausible explanation for this phenomenon lies in the term-based distinctions between text generated by LLMs and human-written content, as evident in Figure 3. On the other hand, the queries are crafted by humans and thus exhibit a style more closely aligned with human-written text.
Neural retrieval models are biased towards LLM-generated texts. An intriguing phenomenon emerges when evaluating neural retrieval models that rely on semantic matching. There is a significant performance gap between their performance on the human-written corpus and LLM-generated corpus. Note that all neural retrieval models achieve a relative performance gain of over $30\%$ when targeting LLM-generated corpus as compared to human-written corpus. These findings suggest an inherent bias in neural retrieval models towards LLM-generated text, which we named the source bias. This bias may arise from the potential exploitation of specific patterns or ‘semantic shortcuts’ embedded in LLM-generated text, given that both neural retrieval models (based on PLMs) and LLMs share a similar Transformer-based (Vaswani et al., 2017) architecture and a similar pretraining paradigm. Moreover, it seems that the information in texts is compressed at the semantic level during LLM generation, resulting in a clearer understanding by neural retrieval models. We will delve deeper into the cause of this bias in the subsequent section.
This set of experiments provides valuable insights into the performance of different retrieval models in the LLM era, underscoring the presence of the source bias within neural retrieval models. Next, we will further uncover more scenarios that exhibit this bias.
Table 4. Performance comparison of different neural retrieval models for mixed human-written and ChatGPT-generated corpora on SciFact+AIGC dataset. The annotations in the table are consistent with Table 3.
Model Target Corpus NDCG@1 NDCG@3 NDCG@5 MAP@1 MAP@3 MAP@5
TF-IDF Human-Written 22.7 36.5 39.5 22.0 32.8 34.6
LLM-Generated 16.7 34.9 37.1 16.0 30.2 31.4
$\text{Relative}~{}~{}\Delta$ 30.5 4.5 6.3 31.6 8.3 9.7
BM25 Human-Written 24.3 38.5 42.7 23.7 34.8 37.3
LLM-Generated 24.3 40.2 42.7 23.1 35.8 37.3
$\text{Relative}~{}~{}\Delta$ 0.0 -4.3 0.0 2.6 -2.8 0.0
ANCE Human-Written 18.0 30.8 33.8 16.5 27.2 29.0
LLM-Generated 24.7 35.6 37.4 24.0 32.7 33.7
$\text{Relative}~{}~{}\Delta$31.4 -14.5 -10.1 -37.0 -18.4 -15.0
BERM Human-Written 16.3 29.9 32.3 14.8 26.0 27.4
LLM-Generated 22.7 32.5 35.3 21.9 29.7 31.4
$\text{Relative}~{}~{}\Delta$32.8 -8.3 -8.9 -38.7 -13.3 -13.6
TAS-B Human-Written 23.0 41.5 44.4 22.2 36.9 38.6
LLM-Generated 28.7 45.5 46.7 27.2 40.9 41.6
$\text{Relative}~{}~{}\Delta$22.1 -9.2 -5.0 -20.2 -10.3 -7.5
Contriever Human-Written 24.0 44.0 47.2 23.3 39.1 41.0
LLM-Generated 33.0 48.3 50.6 31.3 44.0 45.4
$\text{Relative}~{}~{}\Delta$31.6 -9.3 -7.0 -29.3 -11.8 -10.2
Table 5. Bias evaluation of re-ranking models on mixed human-written and Llama2-generated corpora of SciFact+AIGC dataset. The re-ranking methods rerank the top-100 retrieved hits from a first-stage BM25 model. The annotations in the table are consistent with Table 3.
Metrics Target Corpus Base Re-ranking
BM25 +MiniLM +monoT5
NDCG@1 Human-Written 26.7 21.3 19.7
LLM-Generated 21.0 32.7 39.7
$\text{Relative}~{}~{}\Delta$ 23.9 -42.2 -67.3
NDCG@3 Human-Written 40.3 42.8 45.9
LLM-Generated 38.8 47.8 52.9
$\text{Relative}~{}~{}\Delta$ 3.8 -11.0 -14.2
NDCG@5 Human-Written 44.4 46.9 49.0
LLM-Generated 41.5 50.2 54.7
$\text{Relative}~{}~{}\Delta$ 6.8 -6.8 -11.0
MAP@1 Human-Written 25.7 20.8 18.9
LLM-Generated 19.6 30.8 37.8
$\text{Relative}~{}~{}\Delta$ 26.9 -38.8 -66.7
MAP@3 Human-Written 36.7 37.5 39.7
LLM-Generated 34.3 43.6 48.9
$\text{Relative}~{}~{}\Delta$ 6.8 -15.0 -20.8
MAP@5 Human-Written 39.1 40.0 41.6
LLM-Generated 35.9 45.0 50.1
$\text{Relative}~{}~{}\Delta$ 8.5 -11.8 -18.5
3.3. Bias in ChatGPT-Generated Corpus
To strengthen our argument that the source bias is not limited to any specific LLM-generated corpus, we extend our investigation to include ChatGPT, another widely adopted and nearly state-of-the-art LLM. We employ ChatGPT to generate a corpus using the same prompts as those utilized with Llama2 in the above experiments. In Appendix A.4, we present the detailed data quality evaluation of the ChatGPT-generated corpus. Subsequently, in Table 4, we report the evaluation results on the SciFact+AIGC dataset, which contains both human-written and ChatGPT-generated texts. Once again, the results clearly indicate a bias within neural retrieval models, favoring the corpus generated by ChatGPT across all ranking metrics. This observation provides additional substantiation of the presence of source bias within these neural retrieval models.
Table 6. Bias evaluation of re-ranking models on mixed human-written and ChatGPT-generated corpora of SciFact+AIGC dataset. The re-ranking methods rerank the top-100 retrieved hits from a first-stage BM25 model. The annotations in the table are consistent with Table 3.
Metrics Target Corpus Base Re-ranking
BM25 +MiniLM +monoT5
NDCG@1 Human-Written 24.3 18.3 21.3
LLM-Generated 24.3 35.7 39.3
$\text{Relative}~{}~{}\Delta$ 0.0 -64.4 -59.4
NDCG@3 Human-Written 38.5 41.4 46.4
LLM-Generated 40.2 50.1 54.2
$\text{Relative}~{}~{}\Delta$4.3 -19.0 -15.5
NDCG@5 Human-Written 42.7 45.6 48.9
LLM-Generated 42.7 53.0 56.1
$\text{Relative}~{}~{}\Delta$ 0.0 -15.0 -13.7
MAP@1 Human-Written 23.7 17.9 20.5
LLM-Generated 23.1 33.8 37.8
$\text{Relative}~{}~{}\Delta$ 2.6 -61.5 -59.3
MAP@3 Human-Written 34.8 35.8 40.3
LLM-Generated 35.8 45.9 50.0
$\text{Relative}~{}~{}\Delta$2.8 -24.7 -21.5
MAP@5 Human-Written 37.3 38.3 41.7
LLM-Generated 37.3 47.6 51.4
$\text{Relative}~{}~{}\Delta$ 0.0 -21.7 -20.8
3.4. Bias in Re-Ranking Stage
In a typical IR system, there are two primary stages of document filtering. The first stage involves a retriever, responsible for document recall, while the second stage employs a re-ranker, which fine-tunes the ordering of documents within the initially retrieved set. Our investigation has revealed the presence of the source bias in the initial document filtering process (i.e., retrieval stage). Hence, a natural pivotal research question arises: does this bias extend to the ranking models used in the second-stage document filtering process (re-ranking stage), within IR systems?
To delve into this, we choose two representative re-ranking models: MiniLM (Wang et al., 2020) and monoT5 (Nogueira et al., 2020) to rerank the top-100 document list retrieved by a first-stage BM25 model. MiniLM is a cross-encoder model pre-trained on MS MARCO (Nguyen et al., 2016) using a knowledge distillation on three teacher models: BERT-base (Devlin et al., 2019), BERT-large (Devlin et al., 2019), and ALBERT-large (Lan et al., 2019) models. monoT5 is a sequence-to-sequence re-ranker that uses T5 (Raffel et al., 2020) to calculate the relevance score, which is also pretrained on MS MARCO. The results on the SciFact+AIGC dataset with Llama-generated corpus and ChatGPT-generated corpus are presented in Table 5 and Table 6, respectively.
From the results, while even the first-stage retrievers (BM25) may exhibit a preference for human-written content, the second-stage re-rankers once again demonstrate a bias in favor of LLM-generated content. Remarkably, the bias in re-ranking models appears to be more pronounced, as evidenced by the relative percentage difference of $67.3\%$ and $59.4\%$ in NDCG@$1$ for monoT5 in Table 5 and Table 6, respectively. These findings further solidify the pervasiveness of source bias in neural ranking models that rely on PLMs, regardless of the retrieval stage or re-ranking stage.
(a) Relative $\Delta$ on NDCG@$1$
(b) Relative $\Delta$ on MAP@$1$
Figure 6. Performance comparison (NDCG@$1$ and MAP@$1$) of neural retrieval models on SciFact+AIGC dataset under different temperatures of Llama2-generated corpus. A larger negative the Relative $\Delta$ indicates a more severe source bias.
3.5. Bias in Different Temperatures
Temperature is an important hyperparameter used to control the randomness and creativity of text generation in LLMs. A lower temperature sharpens the probability distribution of predicted words, favoring the most likely words. Conversely, a higher temperature flattens the probability distribution, allowing all words to have a more equal chance of being selected. To investigate the impact of temperature on the source bias, we conduct experiments in which Llama2 generated different corpora with varying temperatures (i.e., $T\in\{0.2,1.0\}$). We then examine different neural retrieval models by combining the Llama2-generated corpus with the human-written corpus one by one. The relative gaps in NDCG@$1$ and MAP@$1$ on the SciFact+AIGC dataset are presented in Figure 6. In Appendix A.4, we present detailed data statistics and quality evaluation of the Llama2-generated corpus ($T=1.0$). We also provide the raw results of these retrieval models for the mixed human-written and Llama2-generated corpora ($T=1.0$) in  Table 9 of Appendix  A.5.
The results suggest that all neural retrieval models still exhibit a bias towards LLM-generated text, regardless of the temperature. Furthermore, there is a reduced bias on most neural retrieval models when applied to LLM-generated text with higher temperatures, especially on the re-ranking models. Note that if the temperature coefficient is higher ($T>1.0$), the LLM will not be able to generate coherent meaningful texts, and the generated results will be very random. These findings provide us with potential starting points for analyzing this bias and we will illustrate this in the next section.
4. RQ3: The Cause of Source Bias
In this section, we delve deeper into the cause of the observed source bias. We aim to confirm whether LLM-generated texts, having less noise and more concentrated semantic topics, become easier for neural retrieval models to comprehend. We conduct a series of analyses from the perspective of text compression and provide valuable theoretical insights.
Figure 7. Comparision of the relative singular value (SV) of the different corpus after SVD. The singular values are sorted in descending order from left to right.
Figure 8. Comparision of the relative singular value of LLM-generated corpus with different temperatures. The singular values are sorted in descending order from left to right.
4.1. Viewpoint from Compression
In this analysis, we explore the cause of source bias through a compression perspective, drawing inspiration from recent studies that suggest LLMs act as lossless compressors (Delétang et al., 2023). We postulate that LLMs consistently focus on the major part of the information while filtering out the noise. This, however, does not hold true for humans, as human minds tend to get distracted by chance occurrences. To conduct this examination, we assess the number and strength of topics in both the human-written corpus and the LLM-generated corpus. We employ Singular Value Decomposition (SVD) (Klema and Laub, 1980) on these two corpora. The dimension of the SVD corresponds to the maximum number of topics, and the singular value associated with each topic represents its strength. High singular values predominantly capture primary topic information, whereas low singular values indicate noise. To be more specific, we utilize the OpenAI embedding model to obtain embedding matrices for each corpus within the SciFact+AIGC dataset. Subsequently, we perform SVD on these matrices. The singular values are arranged in descending order, and their comparison to the human-written corpus is visualized in Figure 7.
As illustrated in Figure 7, LLM-generated texts exhibit larger singular values at the top large singular values, while smaller singular values at the tail small singular values. This observation suggests that LLM-generated texts tend to have more focused semantics with less noise, rendering them more suitable for precise semantic similarity calculations. In contrast, human-written texts often contain a wider range of topics and higher levels of noise, making them harder for neural IR models to understand. As a result, this difference in semantic concentration may contribute to the observed bias in neural IR models.
Figure 9. Comparision of the PPL of the different corpora.
Furthermore, we conduct an SVD analysis on LLM-generated documents with different temperature coefficients, as shown in Figure 8. Evidently, for various temperature coefficients of Llama2, the head of the SVD consistently displays higher singular values compared to human-written text. However, the tails of the SVD corresponding to higher temperature coefficients align more closely with the human-written text, while lower temperature coefficients result in tails with smaller singular values. This aligns with our previous findings in Section 3.5, that is, lower temperature results in more conservative and predictable text, and higher temperature results in more creative and diverse text. Thus, lower temperature coefficients lead to more substantial compression, greater topic concentration, and a more pronounced source bias.
4.2. Further Analysis from Perplexity
Considering that most modern neural IR models are grounded on PLMs (Yates et al., 2021; Guo et al., 2020; Zhao et al., 2022), such as BERT (Devlin et al., 2019), Roberta (Liu et al., 2019), and T5 (Raffel et al., 2020), we try to analyze the perplexity of PLMs to further support the conclusion above from the viewpoint of compression that LLM-generated texts can be better understood by PLMs. Perplexity is an important metric for evaluating how well a language model can understand a given text (Azzopardi et al., 2003; Wang et al., 2019). For a specific language model (LM) and a document $d=(d_{0},d_{1},\cdots,d_{S})$ , the log perplexity is defined as the exponentiated average negative log-likelihood of each token in the tokenized sequence of $d$ [^4] For simplicity, we denote the log perplexity as PPL.:
$\text{PPL}(d)=-\frac{1}{S}\left(\sum_{s=1}^{S}\log P_{\text{LM}}(d_{s}|\text{context})\right),$
where $S$ represents the token length of each text $d$ and $P_{\text{LM}}(d_{s})$ is the predicted likelihood of the $s$-th token conditioned on the context. Lower perplexity suggests more confidence and understanding of LM to text patterns, while higher perplexity implies greater uncertainty in predictions, often arising from complex or unpredictable text patterns.
Using the most widely-used LM, BERT (Devlin et al., 2019), as an example, we employ it to calculate the PPL for various corpora. As BERT is not an autoregressive LM, we follow standard practices (Wang et al., 2021; Wang and Cho, 2019) to calculate the predicted likelihood of each token conditioned on the other tokens, i.e.,
$P_{\text{LM}}(d_{s}|\text{context}):=P_{\text{BERT}}(d_{s}|d_{\leq S\backslash\{s\}}).$
The distribution of perplexity for different corpus within the SciFact+AIGC dataset is shown in Figure 9.
Notably, LLM-generated texts consistently exhibit significantly lower perplexity (Figure 9), implying enhanced comprehensibility and higher confidence from BERT’s perspective. Consequently, neural retrieval models based on PLMs can more effectively comprehend LLM-generated texts. This enhanced understanding allows for a more accurate prediction of relevance, leading to the observed source bias in favor of LLM-generated texts. Moreover, as shown in Figure 9, Llama2 with a lower temperature coefficient (i.e., $T=0.2$) generates text with a lower perplexity than the higher temperature coefficient (i.e., $T=1.0$). This observation indicates that LLM with a smaller temperature is less prone to generate unexpected or unusual words, thereby enabling a more effective understanding and grasp of language patterns by the PLMs. To conclude, these observations underscore that texts generated by LLM are more easily comprehensible when processed by PLMs in neural IR models, culminating in the observed source bias.
4.3. Theoretical Analysis and Insights
In  Figure 9, we have compared the PPL for different corpus using the BERT model. In this section, we aim to further provide some theoretical insights into the above observations that LLM-generated texts have a smaller perplexity than human-written texts.
Without loss of generality, we define the PPL in an autoregressive manner. Let $d^{H}$ denote a document written by humans, and $d^{G}$ a document generated by an LLM conditioned on $d^{H}.$ For a given document $d$ and BERT model $\mathcal{B}$, PPL is calculated as
$\text{PPL}(d,\mathcal{B})=-\frac{1}{S}\sum_{s=1}^{S}\log P_{\text{BERT}}(d_{s}|d_{<{s}}).$
Similarly, we use $\text{PPL}(d,\mathcal{H})$ to represent the PPL of document $d$ when evaluated by humans. The PPL of $d^{G}$ conditioned on $d^{H}$ is denoted as
$\text{PPL}(d^{G}\mid d^{H},\mathcal{B})=-\frac{1}{S}\sum_{s=1}^{S}\log P_{\text{BERT}}(d_{s}^{G}|d_{<{s}}^{G},d^{H}).$
When evaluated by humans, we use $\text{PPL}(d^{G}\mid d^{H},\mathcal{H})$ to represent the PPL of $d^{G}$ conditioned on $d^{H}$.
In the theorem below, we introduce two assumptions: Semantic Superiority and Conditional Redundancy, to theoretically establish the sufficient conditions under which $\text{PPL}(d^{G},\mathcal{B})\leq\text{PPL}(d^{H},\mathcal{B})$ holds. Semantic Superiority suggests that the perplexity of human-written texts, when evaluated by humans, is lower than when evaluated by BERT. Conditional Redundancy implies that the perplexity of $d^{G}$, given $d^{H}$, is lower than the perplexity of $d^{H}$ when evaluated directly. This is intuitively true when the information added in generating from $d^{H}$ to $d^{G}$ doesn’t exceed the original information in $d^{H}$. Then we have the following theorem:
Theorem 4.1 ().
Given the following conditions:
  • [leftmargin=0.5cm]
  • Semantic Superiority: human beings outperform BERT in understanding human-written texts, i.e.,
    $\textnormal{PPL}(d^{H},\mathcal{B})-\textnormal{PPL}(d^{H},\mathcal{H})\geq 0.$
  • Conditional Redundancy: generating $d^{G}$ from $d^{H}$ adds less perplexity than $d^{H}$ itself, i.e.,
    $\textnormal{PPL}(d^{H},\mathcal{H})-\textnormal{PPL}(d^{G}\mid d^{H},\mathcal{H})\geq 0.$
  • Bounded Perplexity: there exists a bounded non-negative difference $\epsilon$ in BERT’s perplexity for $d^{G}$ with or without $d^{H},$ i.e.,
    $\textnormal{PPL}(d^{G},\mathcal{B})-\textnormal{PPL}(d^{G}\mid d^{H},\mathcal{B})\leq\epsilon.$
if LLM aligns more closely with BERT than with human when predicting $d^{G}$ given $d^{H}$, such that for any $s\in[S],$
[…]
it follows that
[…]
Proof of Theorem 4.1 can be found in Appendix A.3. In  Theorem 4.1, we assume that there exists an upper bound $\epsilon$ on the increase in perplexity when evaluating $d^{G}$directly, compared to evaluating $d^{G}$ conditioned on $d^{H}$. Moreover, the KL divergence is used to compare the distributions of the document $d^{G}$ conditioned on $d^{H}$ according to the LLM, BERT model, and humans. It is worth to emphasize that inequation  (1) is not the assumption on the understanding capabilities of BERT, LLM, and humans. Instead, this inequation assumes that when predicting $d^{G}$ given $d^{H},$ the predictions by LLM are more closely aligned with those of BERT.
We demonstrate that, when inequation  (1) is satisfied, the perplexity(evaluated by PLMs such as BERT) of $d^{G}$ is lower than that of $d^{H}$. We’d like to emphasize that it is reasonable to expect that inequation  (1) holds true because both LLM and BERT are Transformer-based models that use similar pretraining paradigms. The commonality in model structure and learning paradigms may lead to similar inherent biases in text prediction, making their predictions more aligned with each other.
(a) Results on mixed human-written and Llama2-generated corpora
(b) Results on mixed human-written and ChatGPT-generated corpora
Figure 10. Performance comparison (NDCG@$1$ and MAP@$1$) of neural retrieval models on SciFact+AIGC dataset with different debiased co-efficient $\alpha$ in our proposed debiasing method. A larger positive the Relative $\Delta$ indicates retrieval models rank human-written text more higher than LLM-generated text. The grey dashed line represents Relative $\Delta=0$.
(a) ANCE
(b) BERM
Figure 11. Performance comparison of neural retrievers on only human-written SciFact dataset with different co-efficient $\alpha$ in our proposed debiased method. The results on MAP have a similar tendency and are omitted for simplicity.
5. RQ4: Mitigating Source Bias
Our previous extensive experiments and analysis have revealed that the source bias widely exists in different neural retrievers due to that neural retrievers appear to adopt inherent semantic shortcuts when learning the semantic representation of LLM-generated content. In this section, we propose a simple but effective approach to mitigate source bias by introducing a debiased constraint to the optimization objective. In this way, we can force the neural IR models to focus on modeling semantic relevance rather than the inherent semantic shortcut of the LLM-generated content.
5.1. Our Method: A Debiased Constraint
Our earlier findings of source bias indicate that neural retrieval models tend to rank LLM-generated documents in higher positions. Thus, the motivation of our debiased method is straightforward, which is to force the retrieval models to focus on modeling the semantic relevance and not to assign higher predicted relevance scores to the LLM-generated documents. Specifically, following the practice in Section 2.2, we first generate the corresponding LLM-generated corpus $\mathcal{C}^{G}$ for the original human-written training corpus $\mathcal{C}^{H}$. In this way, we can get the new paired training data $\mathcal{D}=\{(q_{m},d^{H}_{m},d^{G}_{m})\}_{m=1}^{M}$, where each element $(q_{m},d^{H}_{m},d^{G}_{m})$ is a query, human-written document, and LLM-generated document triplet. $d^{H}_{m}$ and $d^{G}_{m}$ are the corresponding human-written and LLM-generated relevant documents for the query $q$, respectively. Then we introduce the debiased constraint, which can be defined as
(2) $\mathcal{L}_{\text{debias}}=\sum_{(q_{m},d^{H}_{m},d^{G}_{m})\in\mathcal{D}}\max\{0,\hat{r}(q,d^{G};\Theta)-\hat{r}(q,d^{H};\Theta)\}$
where the $\hat{r}(q,d^{G};\Theta)$ and $\hat{r}(q,d^{H};\Theta)$ are the predicted relevance score of $(q,d^{G})$ and $(q,d^{H})$ by the retrieval models with parameters $\Theta$, respectively. This constraint can measure and penalize the source bias when the predicted relevance score of $(q,d^{G})$ is greater than that of $(q,d^{H})$. Therefore, we can mitigate the source bias by minimizing this debiased constraint.
Based on the debiased constraint defined in (2), we can define the final loss for training an unbiased neural retrieval model:
(3) $\mathcal{L}=\mathcal{L}_{\text{rank}}+\alpha\mathcal{L}_{\text{debias}}$
where the $\mathcal{L}_{\text{rank}}$ can be any common-used loss for the ranking task, e.g., contrastive loss or regression loss (Zhao et al., 2022; Guo et al., 2020, 2022). And $\alpha$ is the debiased co-efficient that can balance the ranking performance and the degree of the source bias. The larger $\alpha$ indicates the greater penalty on the samples with source bias, leading to the retriever being more likely to rank the human-written text to a higher position.
5.2. Results and Analysis
To evaluate the effectiveness of our proposed debiased method, we equip the debiased constraint defined in Eq. (2) to two representative neural retrieval models, ANCE (Xiong et al., 2020) and BERM (Xu et al., 2023). In the experiments, we vary the debiased co-efficient $\alpha$ within the range of $\{1e\text{-}4,5e\text{-}4,1e\text{-}3,5e\text{-}3,1e\text{-}2\}$. The original retrieval models learned without our proposed debiased constraint are denoted as “w/o debias”. The results for both mixed human-written and Llama2-generated corpora, and mixed human-written and ChatGPT-generated corpora of SciFact+AIGC dataset are presented in Figure 10.
As we can see, as the debiased co-efficient $\alpha$ increases, the Relative $\Delta$ gradually shifts from negative to positive across almost all metrics and mixed datasets. This trend indicates that the neural retrieval models can rank human-written text higher than LLM-generated text with large $\alpha$. This can be attributed to the inclusion of our debiased constraint into the learning objective, which can penalize the biased samples and compel the retrieval models not to assign higher predicted relevance scores to LLM-generated content. Moreover, as shown in Figure 11, our method not only maintains the retrieval performance on the sole human-written corpus but also provides improvements, especially with BERM as the backbone. The reason can be that the introduction of many LLM-generated samples by our method may enhance the retrieval model’s ability to estimate relevance for distinguishing similar samples.
In summary, these empirical results have demonstrated the efficacy of our proposed debiased method in mitigating the source bias to different degrees through the adjustment of the debiased coefficient $\alpha$. This flexibility enables us to customize our debiasing mechanism to meet diverse perspectives and demands. Notably, the decision to maintain equality between the two content sources or favor human-written content can be tailored based on specific requirements and environmental considerations. The optimal strategy for enhancing the IR ecosystem remains an open question that warrants further exploration.
6. Discussion: Sounding the Alarm
In this paper, we have sought to extrapolate the performance of various retrieval models in the LLM era, where the corpus encompasses both human-written and LLM-generated content. Through a rigorous series of experiments and thorough analysis, we have identified that neural retrieval models demonstrate clear preferences for LLM-generated texts, referred to as source bias. This bias, with the burgeoning proliferation of LLMs and AIGC, may raise significant concerns for a variety of aspects.
First, the source bias has the potential to disrupt the ecosystem of information retrieval. On the one hand, our findings underscore the growing dominance of LLM-generated text in IR systems, which makes human-written content gradually unsearchable and could diminish the incentive for content providers to continue creating human-written content. On the other hand, users may find themselves dissatisfied with search results that predominantly consist of LLM-generated content (Cao et al., 2023; Wu et al., 2023a; Chen et al., 2023), leading to a high user churn rate of the IR systems. Consequently, the benefits and experiences of all stakeholders within the IR ecosystem, including platforms, content providers, and users, may erode over time.
Second, the source bias might inadvertently amplify the spread of LLM-generated misinformation. Some recent studies have underscored the ability of LLMs (e.g., ChatGPT) to generate misleading or deceptive information, either unintentionally or intentionally (Su et al., 2023; Chen and Shu, 2023; Pan et al., 2023). Notably, LLM-generated misinformation is proliferating at an alarming pace, outpacing human-generated misinformation, particularly on mainstream news websites (Hanley and Durumeric, 2023; Spitale et al., 2023). What’s more, our findings directly suggest that neural retrieval models tend to favor LLM-generated text, which means that LLM-generated misinformation may occupy higher positions in ranking lists in current IR systems. As a result, the source bias can exacerbate the dissemination of LLM-generated misinformation, potentially posing significant social issues.
Finally, the source bias could be exploited intentionally as a means to manipulate existing retrieval and ranking algorithms. Malicious actors are more likely to leverage LLMs to generate text with a higher likelihood of ranking at a more advantageous position. Indeed, our empirical investigations reveal that malicious users can easily instruct LLMs to generate text with similar semantic information, e.g., through prompts for paraphrase generation and rewriting generation. Thus, this bias could be weaponized as a potential method of attack against neural retrieval models within today’s search engines, reminiscent of earlier web spam link attacks against PageRank  (Gyöngyi et al., 2004).
As discussed above, since LLMs can be readily instructed to generate texts at scale, source bias presents potential tangible and serious threats to the ecosystem of web content, public trust, and online safety. We hope this discussion will raise public awareness and sound the alarm regarding the risks posed by source bias in the LLM era.
7. Related Work
7.1. Large Language Models for IR
The emergence of large language models (LLMs) (Zhao et al., 2023) has ushered in a transformative era across various research domains, such as natural language processing (NLP) (Bubeck et al., 2023; Bang et al., 2023), education (Guo et al., 2023; Nunes et al., 2023), recommender systems (Dai et al., 2023; Fan et al., 2023), finance (Wu et al., 2023b; Huang et al., 2023), and medicine (Alberts et al., 2023; Thirunavukarasu et al., 2023). In the field of IR, much effort has also been made to utilize the remarkable knowledge and capabilities of LLMs to enhance IR systems (Ai et al., 2023; Zhu et al., 2023). In the industry community, an exemplary successful application is New Bing [^5] https://www.bing.com/new, which is an LLM-powered search assistant that adeptly extracts information from various web pages and delivers concise summarized responses to user queries, thereby elevating the overall search experience. In the research community, there has been a proactive exploration of integrating LLMs into the major IR components, including query rewriters (Srinivasan et al., 2022; Wang et al., 2023a), retrievers (Xu et al., 2022; Dai et al., 2022), re-rankers (Sun et al., 2023; Cho et al., 2023), and readers (Shi et al., 2023; Izacard et al., 2022). For a more comprehensive overview of the recent advancements in the LLMs for IR, we recommend referring to the recent survey (Zhu et al., 2023).
7.2. Artificial Intelligence Generated Content
Artificial Intelligence Generated Content (AIGC) is a rapidly advancing field that involves the creation of content using advanced Generative AI (GAI) techniques (Cao et al., 2023; Wu et al., 2023a; Agnese et al., 2020). Unlike traditional content crafted by humans, AIGC can be generated at scale and in considerably less time (Hanley and Durumeric, 2023; Spitale et al., 2023). Recently, the development of LLMs and other GAI models has greatly improved the quality of AIGC content than before. For instance, LLMs such as ChatGPT have shown impressive emergent abilities in generating human-like content (Cao et al., 2023; Wu et al., 2023a). The DALL-E-3 (Betker et al., 2023), another state-of-the-art text-to-image generation system, can follow user instructions to produce high-quality images. Nevertheless, as AIGC becomes more prevalent across myriad domains, ethical concerns and potential risks come into sharper focus (Thakur et al., 2023; Wang et al., 2023b). In fact, inevitably, the GAI models may generate content with bias and discrimination as the large training data always contain bias and toxicity (Zhuo et al., 2023; Deshpande et al., 2023; Bengio et al., 2023). Furthermore, researchers have found that LLMs can be manipulated into generating increasingly deceptive misinformation, posing challenges to online safety (Jiang et al., 2023; Su et al., 2023; Chen and Shu, 2023). In addition, some recent studies indicate that training GAI models with synthetic data could result in the collapse of the next-generation models (Alemohammad et al., 2023; Shumailov et al., 2023; Briesch et al., 2023). Thus, AIGC is a double-edged sword that requires cautious handling.
8. Conclusion and Future Work
In this paper, we provide a preliminary analysis of the impact of the proliferation of generated content on IR systems, which is a pressing and emerging problem in the LLM era. We first introduce two new benchmarks, SciFact+AIGC and NQ320K+AIGC, and build an environment for evaluating IR models in scenarios where the corpus comprises both human-written and LLM-generated texts. Through extensive experiments within this environment, we uncover an unexpected bias of neural retrieval models favoring LLM-generated text. Moreover, we provide an in-depth analysis of this bias from the perspective of text compression. We also introduce a plug-and-play debiased strategy, which shows the potential to mitigate the source bias to different degrees. Finally, we discuss the crucial concerns and potential risks of this bias to the whole web ecosystem.
Our findings provide valuable insights into several promising directions for future research. Firstly, under siege by the LLMs, how to mitigate the observed source bias and ensuring the sustainable development of information ecosystem is a crucial research direction. Secondly, although our focus in this paper centers on information retrieval (IR) systems, the findings and conclusion are not IR-specific. An interesting research question to explore further is whether this bias also manifests in other information systems, such as recommender and advertising systems. Thirdly, it’s essential to recognize that neural models may exhibit biases not only towards generated text but also other data modalities, such as images and multimodal content. Exploring the source bias on tasks beyond texts, such as image-text retrieval, especially in light of emerging generative models like DALL-E-3 (Betker et al., 2023) capable of generating high-quality images, represents a promising research direction.
Appendix A Appendix
A.1. Example of Evaluation Metrics
In this section, we provide a toy example for illustrating the calculation of the evaluation metrics for exploring source bias, as depicted in Figure 12. Note that when calculating metrics for one target corpus, labels of the document from another are masked as negative. For instance, in this toy example, when targeting the human-written corpus, the relevant document $d_{1}^{G}$ generated by LLM is treated as a negative sample. And when calculating the ranking metrics, only consider the rank of the positive human-written documents. When the target corpus is LLM-generated, we adopt the same principle, i.e., only take the rank of the positive LLM-generated documents into account for calculating the ranking metrics.
A.2. Implementation Details
We call the OpenAI API gpt-3.5-turbo-0613 for ChatGPT and llama-2-7b-chat model for LLaMA2  [^6] https://github.com/facebookresearch/llama. If not specified, the temperature of the used LLM for generation corpus is set as 0.2 by default for the experiments. In all our experiments, the nucleus sampling ratio (i.e., top_p) is 1 for both ChatGPT and Llama2.
For the LLM-generated corpus, we conduct post-processing to remove unrelated parts of the original response from LLM like “Sure, here’s a possible rewrite of the text:”. We follow the standard practice in BEIR (Thakur et al., 2021) benchmark to evaluate all the retrieval models. All the models are trained on MS MARCO (Nguyen et al., 2016). For the re-ranking models, we re-rank the top-$100$ documents retrieved by BM25.
A.3. Proof of Theorem 4.1
In this section, we provided the detailed proof of Theorem 4.1
Proof.
Given the following conditions:
  • [leftmargin=0.5cm]
  • Semantic Superiority: human beings outperform BERT in understanding human-authored texts, i.e.,
    $\textnormal{PPL}(d^{H},\mathcal{B})-\textnormal{PPL}(d^{H},\mathcal{H})\geq 0.$
  • Conditional Redundancy: generating $d^{G}$ from $d^{H}$ adds less perplexity than $d^{H}$ itself, i.e.,
    $\textnormal{PPL}(d^{H},\mathcal{H})-\textnormal{PPL}(d^{G}\mid d^{H},\mathcal{H})\geq 0.$
  • Bounded Perplexity: there exists non-negative value $\epsilon$ such that
    $\textnormal{PPL}(d^{G},\mathcal{B})-\textnormal{PPL}(d^{G}\mid d^{H},\mathcal{B})\leq\epsilon.$
Our goal is to prove $\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\left[\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{H},\mathcal{B})\right]\leq 0$ under the condition that LLMs are closer to BERT than to humans when predicting $d^{G}$ given $d^{H}.$ This condition states that, for any $s\in[S],$ the KL divergences satisfy
$D_{\textnormal{KL}}(P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})\|P_{\text{BERT}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})+\epsilon$
$\leq D_{\textnormal{KL}}(P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})\|P_{\text{Human}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})).$
We start by introducing the term $\text{PPL}(d^{H},\mathcal{H}):$
PPL $(d^{G},\mathcal{B})-\text{PPL}(d^{H},\mathcal{B})$
$=\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{H},\mathcal{H})+\text{PPL}(d^{H},\mathcal{H})-\text{PPL}(d^{H},\mathcal{B})$
$\leq\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{H},\mathcal{H}),$
where the last step follows from the Semantic Superiority condition.
PPL $(d^{G},\mathcal{B})-\text{PPL}(d^{H},\mathcal{B})\leq\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{H},\mathcal{H})$
$=\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{L})+\text{PPL}(d^{G}\mid d^{H},\mathcal{L})-\text{PPL}(d^{H},\mathcal{H}).$
Next, we provide upper bounds of $\text{PPL}(d^{G}\mid d^{H},\mathcal{L})-\text{PPL}(d^{H},\mathcal{H}):$
$\text{PPL}(d^{G}\mid d^{H},\mathcal{L})-\text{PPL}(d^{H},\mathcal{H})$
$=\text{PPL}(d^{G}\mid d^{H},\mathcal{L})-\text{PPL}(d^{G}\mid d^{H},\mathcal{H})+\text{PPL}(d^{G}\mid d^{H},\mathcal{H})-\text{PPL}(d^{H},\mathcal{H})$
$\leq\text{PPL}(d^{G}\mid d^{H},\mathcal{L})-\text{PPL}(d^{G}\mid d^{H},\mathcal{H}),$
where the inequality follows from the Conditional Redundancy. Taking expectation on $\text{PPL}(d^{G}\mid d^{H},\mathcal{L})-\text{PPL}(d^{G}\mid d^{H},\mathcal{H}):$
$S\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\left[\text{PPL}(d^{G}\mid d^{H},\mathcal{L})-\text{PPL}(d^{G}\mid d^{H},\mathcal{H})\right]$
$=-\sum_{s=1}^{S}\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\log\frac{P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{L},d^{H})}{P_{\text{Human}}(d_{s}^{G}|d_{<{s}}^{L},d^{H})}$
$=-\sum_{s=1}^{S}\mathbb{E}_{P_{\text{LLM}}(d_{<s}^{G}\mid d^{H})}\mathbb{E}_{P_{\text{LLM}}(d_{s}^{G}\mid d_{<s}^{G},d^{H})}\log\frac{P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{L},d^{H})}{P_{\text{Human}}(d_{s}^{G}|d_{<{s}}^{L},d^{H})}$
$=-\sum_{s=1}^{S}\mathbb{E}_{P_{\text{LLM}}(d_{<s}^{G}\mid d^{H})}\underbrace{D_{\textnormal{KL}}(P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{L},d^{H})\|P_{\text{Human}}(d_{s}^{G}|d_{<{s}}^{L},d^{H}))}_{D_{\textnormal{KL}}(P_{\text{LLM}}\|P_{\text{Human}})}.$
Similarly, $\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{L})$ can be rewritten as:
$\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{L})$
$=\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{B})+\text{PPL}(d^{G}\mid d^{H},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{L}).$
Taking expectation on $\text{PPL}(d^{G}\mid d^{H},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{L}):$
$S\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\left[\text{PPL}(d^{G}\mid d^{H},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{L})\right]$
$=\sum_{s=1}^{S}\mathbb{E}_{P_{\text{LLM}}(d_{<s}^{L}\mid d^{H})}\underbrace{D_{\textnormal{KL}}(P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{L},d^{H})\|P_{\text{BERT}}(d_{s}^{G}|d_{<{s}}^{L},d^{H}))}_{D_{\textnormal{KL}}(P_{\text{LLM}}\|P_{\text{BERT}})}.$
Thus,
$\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\left[\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{H},\mathcal{B})\right]$
$\leq\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\left[\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{B})\right]$
$+\frac{1}{S}\sum_{s=1}^{S}\mathbb{E}_{P_{\text{LLM}}(d_{<s}^{G}\mid d^{H})}(D_{\textnormal{KL}}(P_{\text{LLM}}\|P_{\text{BERT}})-D_{\textnormal{KL}}(P_{\text{LLM}}\|P_{\text{Human}})).$
The final results can be derived by considering the assumptions:
$\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{B})$ $\leq\epsilon$
$D_{\textnormal{KL}}(P_{\text{LLM}}\|P_{\text{BERT}})-D_{\textnormal{KL}}(P_{\text{LLM}}\|P_{\text{Human}})$ $\leq-\epsilon.$
From these, it follows that:
$\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\left[\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{H},\mathcal{B})\right]\leq\epsilon-\epsilon=0.$
Figure 12. A toy example to illustrate how ranking metrics are calculated for each target corpus. In this toy example, given a query, the top 6 documents are retrieved and the rank is from left to right in descending order. Two relevant documents (i.e., $d_{1}^{G}$ and $d_{1}^{H}$) are highlighted with red boxes.
(a) Human vs. ChatGPT (T=0.2)
(b) Human vs. Llama2 (T=0.2)
Figure 13. Semantic embedding visualization of human-written and LLM-generated corpora on SciFact+AIGC datasets.
A.4. More Results of Dataset Quality Validation
In Section 3.2 and 3.5, we also leverage ChatGPT (Temperature=0.2) and Llama2 (Temperature=1.0) to generate LLM-generated corpus for further experiments. For data quality evaluation, we also analyze the semantic similarity. Figure 13 and Figure 14 show the visualization of T-SNE and cosine similarity of the sementic embedding from text-embedding-ada-002777https://platform.openai.com/docs/api-reference/embeddings, respectively. As we can see, the LLM-generated texts almost have the same semantic information of the corresponding original human-written texts. We also conduct experiments with different retrieval models on the sole corpus. As depicted in Table 7 and Table 8, there are no significant performance discrepancies in terms of various ranking metrics between the human-written and LLM-generated corpora, further affirming the quality of our newly constructed datasets.
A.5. Raw Results with Llama2 (T=1.0)
The raw results of retrieval models for the mixed human-written and Llama2-generated corpora ($T=1.0$) are shown in Table 9. Once again, the results highlight the presence of source bias in both neural retrievers and re-rankers.
A.6. Examples of LLM-Generated Corpus
We provide examples of human-written corpus and the corresponding rewritten corpus by Llama2 and ChatGPT on SciFact+AIGC and NQ320K+AIGC dataset, as shown in  subsection A.6 and  subsection A.6, respectively. For different examples in each dataset, we indicate the LLM model and temperature in the sampling strategy.
(a) Llama2 (T=0.2) vs. ChatGPT (T=0.2)
(b) Llama2 (T=0.2) vs. Llama2 (T=1.0)
Figure 14. Distribution of cosine similarity of semantic embedding between LLM-generated and human-written corpora.
Table 7. Performance comparison of different retrieval models on the sole human-written corpus or ChatGPT-generated corpus on SciFact+AIGC dataset. The temperature of ChatGPT is set as 0.2.
Model Type Model Corpus NDCG@1 NDCG@3 NDCG@5 MAP@1 MAP@3 MAP@5
Lexical TF-IDF Human-Written 42.0 49.5 52.7 40.7 47.1 49.0
LLM-Generated 41.7 49.9 52.1 39.4 47.1 48.6
BM25 Human-Written 46.0 54.2 56.3 43.8 51.5 52.8
LLM-Generated 46.7 53.9 56.0 44.4 51.5 52.7
Neural ANCE Human-Written 38.7 44.3 46.5 36.3 41.9 43.3
LLM-Generated 40.3 45.5 47.9 38.2 43.3 44.9
BERM Human-Written 37.0 42.1 44.2 34.7 39.7 41.3
LLM-Generated 37.3 43.2 45.0 35.1 40.8 42.0
TAS-B Human-Written 52.7 58.1 60.2 49.9 55.6 57.2
LLM-Generated 50.0 57.3 59.8 48.2 54.9 56.5
Contriever Human-Written 54.0 61.8 63.2 51.4 58.9 60.0
LLM-Generated 55.7 62.1 64.8 53.3 59.7 61.5
Table 8. Performance comparison of different retrieval models on the sole human-written corpus or Llama2-generated corpus on SciFact+AIGC datasets. The temperature of Llama2 is set as 1.0.
Model Type Model Corpus NDCG@1 NDCG@3 NDCG@5 MAP@1 MAP@3 MAP@5
Lexical TF-IDF Human-Written 42.0 49.5 52.7 40.7 47.1 49.0
LLM-Generated 42.0 48.9 50.7 39.8 46.4 47.5
BM25 Human-Written 46.0 54.2 56.3 43.8 51.5 52.8
LLM-Generated 46.0 52.4 54.7 43.6 50.0 51.5
Neural ANCE Human-Written 38.7 44.3 46.5 36.3 41.9 43.3
LLM-Generated 39.7 44.9 46.5 36.9 42.6 43.7
BERM Human-Written 37.0 42.1 44.2 34.7 39.7 41.3
LLM-Generated 38.0 42.3 43.5 35.4 40.0 41.0
TAS-B Human-Written 52.7 58.1 60.2 49.9 55.6 57.2
LLM-Generated 52.3 55.5 59.0 49.5 53.7 55.9
Contriever Human-Written 54.0 61.8 63.2 51.4 58.9 60.0
LLM-Generated 52.0 59.0 62.0 49.6 56.4 58.4
Table 9. Performance comparison of different retrieval models for mixed human-written and Llama2-generated corpora on SciFact+AIGC dataset. The temperature of Llama2 is set as 1.0.
Model Type Model Corpus NDCG@1 NDCG@3 NDCG@5 MAP@1 MAP@3 MAP@5
Lexical TF-IDF Human-Written 21.7 36.5 39.3 20.9 32.6 34.3
LLM-Generated 15.3 32.8 35.1 14.5 28.2 29.6
$\text{Relative}~{}~{}\Delta$ 34.6 10.7 11.3 36.2 14.5 14.7
BM25 Human-Written 24.3 39.7 43.2 23.2 35.6 37.6
LLM-Generated 22.7 39.4 41.6 21.4 35.0 36.3
$\text{Relative}~{}~{}\Delta$ 6.8 0.8 3.8 8.1 1.7 3.5
Neural ANCE Human-Written 15.7 30.8 33.1 14.1 26.7 28.1
LLM-Generated 23.7 34.2 36.6 22.6 31.2 32.6
$\text{Relative}~{}~{}\Delta$40.6 -10.5 -10.0 -46.3 -15.5 -14.8
BERM Human-Written 16.7 29.6 31.8 15.9 26.3 27.6
LLM-Generated 22.3 32.1 33.9 20.4 29.0 30.1
$\text{Relative}~{}~{}\Delta$28.7 -8.1 -6.4 -24.8 -9.8 -8.7
TAS-B Human-Written 21.0 41.0 43.2 20.0 35.9 37.2
LLM-Generated 32.3 44.5 46.8 30.7 41.1 42.4
$\text{Relative}~{}~{}\Delta$42.4 -8.2 -8.0 -42.2 -13.5 -13.1
Contriever Human-Written 25.0 43.9 47.8 24.5 39.3 41.6
LLM-Generated 29.0 44.4 47.4 27.4 40.2 41.9
$\text{Relative}~{}~{}\Delta$14.8 -1.1 0.8 -11.2 -2.3 -0.7
Re-ranking BM25+MiniLM Human-Written 22.0 42.0 47.0 21.5 37.3 39.8
LLM-Generated 31.0 47.0 50.0 29.3 42.2 43.9
$\text{Relative}~{}~{}\Delta$34.0 -11.2 -6.2 -30.7 -12.3 -9.8
BM25+monoT5 Human-Written 24.7 47.1 50.5 23.8 41.5 43.6
LLM-Generated 35.0 51.6 53.8 32.9 47.0 48.4
$\text{Relative}~{}~{}\Delta$34.5 -9.1 -6.3 -32.1 -12.4 -10.4
Table 10. Examples of human-written and LLM-generated texts on SciFact+AIGC dataset.
Human
Likelihood ratios are one of the best measures of diagnostic accuracy, although they are seldom used, because interpreting them requires a calculator to convert back and forth between “probability” and “odds” of disease. This article describes a simpler method of interpreting likelihood ratios, one that avoids calculators, nomograms, and conversions to “odds” of disease. Several examples illustrate how the clinician can use this method to refine diagnostic decisions at the bedside.
[…]
Llama2 (Temperature=0.2)
Likelihood ratios are a valuable tool for assessing the accuracy of diagnostic tests, yet they are often overlooked due to the complexity of interpreting them. Converting between "probability" and "odds" of disease requires a calculator, which can be a barrier for clinicians at the bedside. This article proposes a simpler method of interpreting likelihood ratios, eliminating the need for calculators, nomograms, or conversions to "odds" of disease. With this approach, clinicians can make more informed diagnostic decisions in real-time. Several examples demonstrate how this method can be applied in practice.
[…]
Llama2 (Temperature=1.0)
Likelihood ratios are a powerful tool for evaluating the accuracy of diagnostic tests, yet they are often overlooked due to the complexity of interpreting them. Conversion between "probability" and "odds" of disease requires a calculator, which can make the process challenging for clinicians. This article presents a simplified approach to interpreting likelihood ratios, eliminating the need for calculators, nomograms, or conversions to "odds" of disease. By following this method, clinicians can improve their diagnostic decision-making at the bedside through practical examples.
[…]
ChatGPT (Temperature=0.2)
The use of likelihood ratios is considered one of the most effective ways to assess diagnostic accuracy. However, they are not commonly utilized due to the complexity of interpreting them, which involves converting between the "probability" and "odds" of disease using a calculator. This article presents a simpler approach to interpreting likelihood ratios that eliminates the need for calculators, nomograms, and conversions to disease odds. Through various examples, it demonstrates how clinicians can employ this method to enhance diagnostic decision-making directly at the patientś bedside.
Table 11. Examples of human-written and LLM-generated texts on NQ320K+AIGC dataset.
Case 1
[…]
2018 FIFA World Cup disciplinary record In the 2018 FIFA World Cup , which took place in Russia from 14 June to 15 July 2018 , the main disciplinary actions taken against players came in the form of red and yellow cards . A player sanctioned with a red card was sent off from the pitch and could not be replaced . Furthermore , the player was automatically banned from his team ’s next match . After a straight red card , FIFA conducted a hearing and considered extending this ban beyond one match . If the ban extended beyond the end of the World Cup finals ( for example , a player was sent off in his team ’s last match ) , it had to be served in the team ’s next competitive international match ( es ) . A player also received a one match ban if he picked up a single yellow card in two different games from the beginning of the group stage until the end of the quarter - finals , where single yellow cards were negated . However , such a ban did not carry over beyond the World Cup finals if the
[…]
Llama2 (Temperature=0.2)
In the 2018 FIFA World Cup, which took place in Russia from June 14 to July 15, 2018, the main disciplinary actions against players were red and yellow cards. A player sent off with a red card was unable to be replaced, and was automatically banned from their team’s next match. Following a straight red card, FIFA conducted a hearing and could have extended this ban beyond one match. If the ban extended beyond the end of the World Cup finals (e.g., a player was sent off in their team’s last match), it had to be served in their team’s next competitive international match. A player received a one-match ban if they accumulated a single yellow card in two or more games from the start of the group stage to the end of the quarter-finals, where single yellow cards were voided. However, this ban did not carry over beyond the World Cup finals if the player’s team was eliminated before the quarter-finals.
Case 2
[…]
San Andreas fault The San Andreas Fault is a continental transform fault that extends roughly 1,200 kilometers ( 750 mi ) through California . It forms the tectonic boundary between the Pacific Plate and the North American Plate , and its motion is right - lateral strike - slip ( horizontal ) . The fault divides into three segments , each with different characteristics and a different degree of earthquake risk , the most significant being the southern segment , which passes within about 35 miles ( 56 km ) of Los Angeles . The slip rate along the fault ranges from 20 to 35 mm ( 0.79 to 1.38 in ) / yr . The fault was first identified in 1895 by Professor Andrew Lawson of UC Berkeley , who discovered the northern zone . It is often described as having been named after San Andreas Lake , a small body of water that was formed in a valley between the two plates . However , according to some of his reports from 1895 and 1908 , Lawson actually named it after the surrounding San Andreas Valley . Following the 1906 San Francisco earthquake , Lawson concluded that the
[…]
Llama2 (Temperature=0.2)
The San Andreas Fault is a continental transform fault that spans approximately 1,200 kilometers (750 miles) through California, serving as the tectonic boundary between the Pacific Plate and the North American Plate. The fault’s motion is right-lateral strike-slip (horizontal), causing the plates to move horizontally past each other. The fault is divided into three segments, each with distinct characteristics and varying degrees of earthquake risk, with the southern segment posing the greatest risk, passing within approximately 35 miles (56 km) of Los Angeles. The slip rate along the fault ranges from 20 to 35 mm (0.79 to 1.38 in) per year. The fault was first identified in 1895 by Professor Andrew Lawson of UC Berkeley, who discovered the northern zone. Contrary to popular belief, the fault was not named after San Andreas Lake, a small body of water located in a valley between the two plates. Instead, Lawson actually named it after the surrounding San Andreas Valley, as revealed in his reports from 1895 and 1908. Following the 1906 San Francisco earthquake, Lawson concluded that the fault was responsible for the devastating event.

### v3

@@
-LLMs may Dominate Information Access: Neural Retrievers are Biased Towards LLM-Generated Texts
+Neural Retrievers are Biased Towards LLM-Generated Content
 Abstract.
-Recently, the emergence of large language models (LLMs) has revolutionized the paradigm of information retrieval (IR) applications, especially in web search.
-With their remarkable capabilities in generating human-like texts, LLMs have created enormous texts on the Internet.
-As a result, IR systems in the LLMs era are facing a new challenge: the indexed documents now are not only written by human beings but also automatically generated by the LLMs.
+Recently, the emergence of large language models (LLMs) has revolutionized the paradigm of information retrieval (IR) applications, especially in web search, by generating vast amounts of human-like texts on the Internet.
+As a result, IR systems in the LLM era are facing a new challenge: the indexed documents are now not only written by human beings but also automatically generated by the LLMs.
 How these LLM-generated documents influence the IR systems is a pressing and still unexplored question.
-In this work, we conduct a quantitative evaluation of different IR models in scenarios where both human-written and LLM-generated texts are involved.
+In this work, we conduct a quantitative evaluation of IR models in scenarios where both human-written and LLM-generated texts are involved.
 Surprisingly, our findings indicate that neural retrieval models tend to rank LLM-generated documents higher.
-We refer to this category of biases in neural retrieval models towards the LLM-generated text as the source bias.
+We refer to this category of biases in neural retrievers towards the LLM-generated content as the source bias.
 Moreover, we discover that this bias is not confined to the first-stage neural retrievers, but extends to the second-stage neural re-rankers.
-Then, we provide an in-depth analysis from the perspective of text compression and observe that neural models can better understand the semantic information of LLM-generated text, which is further substantiated by our theoretical analysis.
-To mitigate the source bias, we also propose a plug-and-play debiased constraint for the optimization objective, and experimental results show the effectiveness.
+Then, in-depth analyses from the perspective of text compression indicate that LLM-generated texts exhibit more focused semantics with less noise, making it easier for neural retrieval models to semantic match.
+To mitigate the source bias, we also propose a plug-and-play debiased constraint for the optimization objective, and experimental results show its effectiveness.
 Finally, we discuss the potential severe concerns stemming from the observed source bias and hope our findings can serve as a critical wake-up call to the IR community and beyond.
-To facilitate future explorations of IR in the LLM era, the constructed two new benchmarks and codes will later be available at https://github.com/KID-22/LLM4IR-Bias.
-copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: XXXX; August 03–05, 2023; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
+To facilitate future explorations of IR in the LLM era, the constructed two new benchmarks are available at https://github.com/KID-22/Source-Bias.
+Source Bias, Information Retrieval, LLM-Generated Texts, Artificial Intelligence Generated Content
+journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671882isbn: 979-8-4007-0490-1/24/08ccs: Information systems Information retrieval
 1. Introduction
@@
 This influx of LLM-generated content has fundamentally reshaped the digital ecosystem, challenging conventional paradigms of content creation, dissemination, and information access on the Internet (Ai et al., 2023; Zhu et al., 2023).
-Meanwhile, information retrieval (IR) systems have become the keystone of how individuals navigate and access information within the vast expanse of the Internet (Singhal et al., 2001; Manning, 2009).
-The core function of an IR system is to efficiently retrieve relevant documents from the target corpus in response to users’ queries (Liu et al., 2009; Li, 2022; Xu and Li, 2007).
-As illustrated in Figure 1, in the era preceding the widespread emergence of LLMs, IR systems primarily revolved around the retrieval of relevant documents solely from the human-written corpus.
-However, with the rise of LLMs, IR systems are confronted with a profound paradigm shift.
-That is, as LLMs have significantly driven the proliferation of AIGC, the corpus of IR systems now gradually includes both human-written and LLM-generated texts.
+Meanwhile, information retrieval (IR) systems have become indispensable for navigating and accessing the Internet’s vast information landscape (Singhal et al., 2001; Manning, 2009).
+As illustrated in Figure 1, in the era preceding the widespread emergence of LLMs, IR systems focused on retrieving documents solely from the human-written corpus in response to users’ queries (Liu et al., 2009; Li, 2022; Xu and Li, 2007).
+However, the proliferation of AIGC driven by LLMs has expanded the corpus of IR systems to include both human-written and LLM-generated texts.
 Consequently, this paradigm shift raises a fundamental research question: What is the impact of the proliferation of generated content on IR systems?
-In this LLM era, we aim to explore whether existing retrieval models tend to prioritize LLM-generated text over human-written text, even when both convey similar semantic information.
+We aim to explore whether existing retrieval models tend to prioritize LLM-generated text over human-written text, even when both convey similar semantic information.
 If this holds, LLMs may dominate information access, particularly as their generated content is rapidly growing on the Internet (Hanley and Durumeric, 2023).
-To approach the fundamental research question, we decompose it into three specific research questions.
-The first question is RQ1: How to construct a simulation environment to evaluate IR models in the LLM era?
-As the LLM-generated content is currently unidentifiable on the Internet, there are no public retrieval benchmarks where the corpus contains both human-written text and LLM-generated text.
-To construct a more realistic evaluation environment, we introduce a natural approach to synthesize LLM-generated data without incurring significant costs for human annotation of relevant labels.
+To approach the fundamental research question, we decompose it into four specific research questions.
+The first question is RQ1: How to construct an environment to evaluate IR models in the LLM era?
+Given the lack of public retrieval benchmarks encompassing both human-written and LLM-generated texts, we propose an innovative and practical method to create such a realistic evaluation environment without the need of costly human annotation.
 Specifically, we leverage the original human-written texts as the instruction conditions to prompt LLMs to generate rewritten text copies while preserving the same semantic meaning.
-In this way, we can confidently assign the relevant labels for LLM-generated data.
-Extensive empirical analysis validates the quality of our constructed environment, as it can accurately reflect real-world situations of IR in the LLM era.
+In this way, we can confidently assign the same relevancy labels to LLM-generated data according to the original labels.
+Extensive empirical analysis validates the quality of our constructed environment, demonstrating its effectiveness in mirroring real-world IR scenarios in the LLM era.
 As a result, we introduce two new benchmarks, SciFact+AIGC and NQ320K+AIGC, tailored for IR research in the LLM era.
-Based on the constructed environment, we can explore further RQ2: Are and where are retrieval models biased towards LLM-generated texts?
-We conduct systematical experiments with various representative and state-of-the-art retrieval models, ranging from traditional lexical models to modern neural retrieval models (Zhao et al., 2022; Yates et al., 2021; Guo et al., 2022).
-Surprisingly, our findings indicate that neural retrieval models are biased towards LLM-generated texts, i.e., rank LLM-generated texts in a higher position.
-We refer to this as source bias, as the neural retrievers favor texts from specific sources (i.e., LLM-generated).
-Our further experiments indicate that the source bias not only extends to the second-stage neural re-rankers but also manifests in a more severe manner.
-These findings further corroborate the prevalence of source bias in neural ranking models that rely on pretrained language models (PLMs) (Yates et al., 2021; Guo et al., 2020; Zhao et al., 2022).
-In other words, these neural ranking models appear to adopt some inherent semantic shortcuts to learn the semantic representation of documents generated by LLM.
+Figure 2. The overall paradigm of the proposed evaluation framework for IR in the LLM era.
+With the constructed environment, we further explore RQ2: Are retrieval models biased towards LLM-generated content?
+We conduct comprehensive experiments with various representative retrieval models, ranging from traditional lexical models to modern neural models based on pretrained language models (PLMs) (Guo et al., 2020; Zhao et al., 2023a; Yates et al., 2021; Guo et al., 2022).
+Surprisingly, we uncover that neural retrievers are biased towards LLM-generated texts, i.e., tend to rank LLM-generated texts in higher positions.
+We refer to this as source bias, as the neural retrievers favor content from specific sources (i.e., LLM-generated content).
+Further experiments indicate that the source bias not only extends to the second-stage neural re-rankers from the first-stage retrieval but also manifests more severely.
+These findings corroborate the prevalence of source bias in neural retrieval models.
 Then, what we are curious about is RQ3: Why are neural retrieval models biased towards LLM-generated texts?
-Inspired by the recent studies that posit LLMs are lossless compressors (Delétang et al., 2023), we aim to analyze the cause of source bias from the viewpoint of text compression.
-We first examine the singular value (Klema and Laub, 1980) of different corpora and uncover that LLM-generated text typically maintains more focused semantics with minimal noise, rendering it better understood for precise semantic matching.
-Furthermore, to better examine whether the neural retrieval model based on PLMs can better understand the LLM-generated texts, we conduct an in-depth analysis for perplexity.
-Our experimental and theoretical analysis suggest that LLM-generated text consistently yields lower perplexity scores, signifying greater comprehensibility and higher confidence from the perspective of PLMs.
-These observations confirm that LLM-generated text is more readily understandable to neural retrieval models, thereby resulting in source bias.
-As our analysis has revealed source bias, we aim to further study the RQ4: How can we mitigate the source bias in neural retrieval models?
-Specifically, we propose a simple but effective debiased constraint to penalize biased samples during optimization.
-In this way, we can force the retrieval models to focus on modeling semantic relevance rather than the inherent shortcut between neural retrievers and LLM-generated content, leading to the mitigation of source bias.
-This debiased constraint can be plugged and played into the ranking optimization objective of various neural retrieval models.
-Besides, our approach offers the capability to control the degree of bias removal, providing flexibility in determining whether the two sources of content should be treated equally or biased toward human-written content according to specific requirements and environmental considerations.
-Finally, we discuss the potential emerging concerns stemming from source bias.
-First, the presence of source bias poses a significant risk of gradually rendering human-written content less accessible, potentially causing a disruption in the content ecosystem.
-This concern is further compounded by the growing prevalence of LLM-generated text on the Internet(Hanley and Durumeric, 2023; Bengio et al., 2023).
-Second, there is the risk that source bias may amplify the spread of misinformation, especially considering the potential of LLMs to generate deceptive content, whether intentionally or not (Chen and Shu, 2023; Su et al., 2023).
-Third, source bias may be maliciously exploited to manipulate existing retrieval and ranking algorithms, creating a precarious vulnerability that could be weaponized by malicious actors.
+Inspired by the recent studies positing LLMs as lossless compressors (Delétang et al., 2023), we analyze the cause of source bias from the viewpoint of text compression.
+Our analysis of singular values (Klema and Laub, 1980) in different corpora reveals that LLM-generated texts exhibit more focused semantics with minimal noise, enhancing their suitability for semantic matching.
+Furthermore, our in-depth perplexity analysis shows that LLM-generated texts consistently achieve lower perplexity scores, which indicates a higher degree of comprehensibility and confidence from the PLM’s perspective.
+These observations collectively suggest that LLM-generated texts are more readily understandable to PLM-based neural retrievers, thereby resulting in source bias.
+Finally, we try to answer RQ4: How to mitigate source bias in neural retrieval models?
+To tackle this, we propose an intuitive yet effective debiased constraint.
+This constraint is designed to penalize biased samples during the optimization process, thereby shifting the focus of retrieval models from exploiting inherent shortcuts to emphasizing semantic relevance.
+Besides, our debiased constraint is model-agnostic and can be plugged and played to the ranking optimization objectives of various neural retrieval models.
+Furthermore, it offers the capability to control the degree of bias removal, offering the flexibility to balance the treatment between the two sources of content based on specific requirements and environmental considerations.
+Last but not least, we discuss the potential emerging concerns stemming from source bias, highlighting the risk of human-written content being gradually inaccessible, especially due to the rapidly increasing LLM-generated content on the Internet (Hanley and Durumeric, 2023; Bengio et al., 2023).
+Furthermore, source bias could be maliciously exploited to manipulate algorithms and potentially amplify the spread of misinformation, posing a threat to online security.
 In light of these pressing issues, we hope that our findings serve as a resounding wake-up call to all stakeholders involved in IR systems and beyond.
 In summary, the contributions of this paper are as follows:
-(1) This is the first work that introduces the paradigm of IR systems in the LLM era, where the retrieval corpus consists of both human-written texts and LLM-generated texts.
-This scenario is a more realistic setting, considering the growing prosperity of AIGC.
-(2) We uncover the source bias in the LLM era, a new inherent bias in neural retrieval models, where both the neural retrievers and re-rankers rank the LLM-generated text in a higher position than human-written text with the almost same semantic relevance.
-(3) We provide an in-depth analysis and theoretical insights of source bias from the viewpoint of text compression.
-The analysis indicates that LLM-generated text maintains more focused semantics with minimal noise and is more readily comprehensible for neural retrieval models.
-(4) We propose a debiased constraint to penalize the bias of neural retrieval models during the optimization and experimental results demonstrate the effectiveness in mitigating source bias.
-(5) We also provide two new benchmarks, SciFact+AIGC and NQ320K+AIGC, which contain both high-quality human-written and various LLM-generated corpus and corresponding relevant labels.
+(1) We introduce a more realistic paradigm of IR systems considering the growing prosperity of AIGC, where the retrieval corpus consists of both human-written and LLM-generated texts.
+We then uncover a new inherent bias in both neural retrieves and re-rankers preferring LLM-generated content, termed as source bias.
+(2) We provide an in-depth analysis and insights of source bias from a text compression perspective, which indicates that LLM-generated texts maintain more focused semantics with minimal noise and are more readily comprehensible for neural retrievers.
+(3) We propose a debiased constraint to penalize the biased samples during optimization, and experimental results demonstrate its effectiveness in mitigating source bias in different degrees.
+(4) We also provide two new benchmarks, SciFact+AIGC and NQ320K+AIGC, which contain both high-quality human-written and various LLM-generated corpus and corresponding relevancy labels.
 We believe these two benchmarks can serve as valuable resources for facilitating future research of IR in the LLM era.
-Figure 2. The overall paradigm of the proposed evaluation framework for IR in the LLM era.
 2. RQ1: Environment Construction
-With the increasing usage of LLMs in generating texts (e.g., paraphrasing or rewriting texts with LLMs), the corpus of IR systems includes both human-written and LLM-generated texts nowadays.
-In this section, we introduce a novel and natural framework for the quantitative evaluation of retrieval models in the LLM era, i.e., in scenarios where the corpus includes both human-written and LLM-generated texts, as shown in Figure 2.
-To better align with the real-world scenarios in the LLM era, the simulated environments should meet the following two essential criteria.
-Firstly, it is imperative to distinguish between human-written texts and LLM-generated texts within the corpus.
-Secondly, we need access to pertinent labels for LLM-generated data in response to queries.
-To ensure the most effective evaluation of retrieval model preferences, every piece of human-written text should have a corresponding LLM-generated counterpart with consistent content.
-2.1. Preliminary
-2.1.1. Notation
-Formally, in the Pre-LLM era, given a query text $q\in\mathcal{Q}$ where $\mathcal{Q}$ is the set of all queries, the traditional IR system aims to retrieve a list of top-$K$ relevant documents $\{d^{(1)},d^{(2)},\ldots,d^{(K)}\}$ from a corpus $\mathcal{C}^{H}=\{d_{1}^{H},d_{2}^{H},\ldots d_{N}^{H}\}$ which consists of $N$ human-written documents.
+With the increasing usage of LLMs in generating texts (e.g., paraphrasing or rewriting), the corpus of IR systems includes both human-written and LLM-generated texts nowadays.
+Constructing an IR dataset in the LLM era typically involves two steps: collecting both human-written and LLM-generated corpora and then employing human evaluators to annotate relevancy labels for each query-document pair.
+Given that LLM-generated content is currently unidentifiable (Sadasivan et al., 2023) and the significant cost of human annotation, we introduce a natural and practical framework for quantitatively evaluating retrieval models in the LLM era, as shown in Figure 2.
+To better align with real-world scenarios, the evaluation environments should meet the following three essential criteria.
+Firstly, it is imperative to distinguish between human-written and LLM-generated texts within the corpus.
+Secondly, we need access to relevancy labels for LLM-generated data in response to queries.
+Thirdly, each human-written text should better have a corresponding LLM-generated counterpart with the same semantics, ensuring the most effective and fair evaluation.
+Table 1. Statistics of the constructed two datasets. Avg. Doc / Query means the average number of relevant documents per query.
+Dataset # Test Queries # Avg. Query Length Human-Written Corpus Llama2-Generated Corpus ; ChatGPT-Generated Corpus
+# Corpus Avg. Doc Length Avg. Doc / Query # Corpus Avg. Doc Length Avg. Doc / Query
+SciFact+AIGC 300 12.38 5,183 201.81 1.1 5,183 ; 5,183 192.66 ; 203.57 1.1 ; 1.1
+NQ320K+AIGC 7,830 9.24 109,739 199.79 1.0 109,739 ; – 174.49 ; – 1.0 ; –
+2.1. Notation
+Formally, in the Pre-LLM era, given a query $q\in\mathcal{Q}$ where $\mathcal{Q}$ is the set of all queries, the traditional IR system aims to retrieve a list of top-$K$ relevant documents $\{d^{(1)},d^{(2)},\ldots,d^{(K)}\}$ from a corpus $\mathcal{C}^{H}=\{d_{1}^{H},d_{2}^{H},\ldots d_{N}^{H}\}$ which consists of $N$ human-written documents.
 However, in the era of LLMs, there is also LLM-generated text in the corpus.
-To simulate and evaluate the IR models in the LLM era, we also create an additional corpus $\mathcal{C}^{G}=\{d_{1}^{G},d_{2}^{G},\ldots,d_{N}^{G}\}$ where each document is generated by a LLM, e.g., $d_{1}^{G}$ can be created by ChatGPT by constructing a prompt that asks ChatGPT to rewrite $d_{1}^{H}$ while preserving its original meaning.
-Consequently, given a user query $q$, the primary objective of a retriever in the LLM era is to return the top-$K$ relevant documents from the combined corpus $\mathcal{C}=\mathcal{C}^{H}\bigcup\mathcal{C}^{G}$.
-2.1.2. Retrieval Models
-IR models have witnessed a significant evolution, shifting from traditional lexical approaches primarily reliant on term matching to the recent neural IR models founded on semantic matching (Zhao et al., 2022; Guo et al., 2022; Yates et al., 2021; Gonzalo et al., 2014).
-In this study, we select the following representative retrieval models to conduct a thorough exploration of their performance in scenarios encompassing both human-written and LLM-generated texts:
-Lexical Retrieval Models are based on the exact term matching between high-dimensional sparse vectors of query and document.
-We choose the widely used two approaches TF-IDF (Sparck Jones, 1972) and BM25 (Robertson et al., 2009) in our experiments.
-Neural Retrieval Models utilize neural networks to learn semantic relationships between query and document to improve the lexical gap (Li et al., 2014; Zhao et al., 2022; Yates et al., 2021; Guo et al., 2020).
-We select the following four representative neural retrieval models: ANCE (Xiong et al., 2020) is a dense retrieval model that selects hard training negatives globally from the entire corpus, using an asynchronously updated ANN index.
-BERM (Xu et al., 2023) is also a dense retrieval model that focuses on learning the domain-invariant features in information retrieval tasks.
-TAS-B (Hofstätter et al., 2021) is a bi-encoder trained with a balanced margin sampling technique that samples queries out of a cluster per batch, ensuring efficiency in the sampling process.
-Contriever (Izacard et al., 2021) uses momentum contrastive learning to train dense retrievers from text paragraphs and obtain state-of-the-art zero-shot retrieval performance.
-Table 1. Statistics of the constructed two datasets. Avg. Doc / Query means the average number of relevant documents per query.
-Dataset # Test Queries # Avg. Query Length Human-Written Corpus Llama2-Generated/ChatGPT-Generated Corpus
-# Corpus Avg. Doc Length Avg. Doc / Query # Corpus Avg. Doc Length Avg. Doc / Query
-SciFact+AIGC 300 12.38 5,183 201.81 1.1 5,183/5,183 192.66/203.57 1.1/1.1
-NQ320K+AIGC 7,830 9.24 109,739 199.79 1.0 109,739/- 174.49/- 1.0/-
+To evaluate the IR models in the LLM era, we also create an additional LLM-generated corpus $\mathcal{C}^{G}=\{d_{1}^{G},d_{2}^{G},\ldots,d_{N}^{G}\}$ where each document is generated by a LLM, e.g., $d_{1}^{G}$ can be created by prompting ChatGPT to rewrite $d_{1}^{H}$ while preserving its original semantics information.
+Consequently, given a query $q$, the objective of a retriever in the LLM era is to return the top-$K$ relevant documents from the mixed corpus $\mathcal{C}=\mathcal{C}^{H}\bigcup\mathcal{C}^{G}$.
 2.2. Constructing IR Datasets in the LLM Era
-To better align with real-world IR scenarios in the LLM era, we first prompt LLMs to rewrite human-written corpus to build two new standard retrieval datasets: SciFact+AIGC and NQ320K+AIGC.
+In this section, we prompt LLMs to rewrite human-written corpus to build two new standard retrieval datasets: SciFact+AIGC and NQ320K+AIGC.
 These two new datasets can serve as valuable resources to facilitate future research of IR in the LLM era.
-(a) Jaccard Similarity
-(b) Overlap
-Figure 3. Distribution of term Jaccard similarity and overlap between Llama2-generated and human-written corpora.
-(a) SciFact+AIGC
-(b) NQ320K+AIGC
-Figure 4. Semantic embedding visualization of different corpora on SciFact+AIGC and NQ320K+AIGC datasets.
-2.2.1. Human-Written Corpus
-We first choose two widely used retrieval datasets written by humans in the Pre-LLM era as the seed data: SciFact and NQ320K.
-SciFact [^1] https://allenai.org/data/scifact (Wadden et al., 2020) dataset aims to retrieve evidence from the research literature containing scientific paper abstracts for fact-checking.
-NQ320K [^2] https://ai.google.com/research/NaturalQuestions (Kwiatkowski et al., 2019) is based on the Natural Questions (NQ) dataset from Google, where the documents are gathered from Wikipedia pages, and the queries are natural language questions.
-Following the practice in BEIR benchmark (Thakur et al., 2021), we process these two datasets in a standard format: corpus $\mathcal{C}^{H}$, queries $\mathcal{Q}$, and labels $\mathcal{R}^{H}=\{(q_{m},d^{H}_{m},r_{m})\}^{M}_{m=1}$, where $M$ is the number of labeled query-document pairs in the corpus.
-2.2.2. LLM-Generated Corpus
-Constructing an IR dataset in the LLM era typically involves two steps: collecting the LLM-generated corpus $\mathcal{C}^{G}$ and employing human evaluators from crowdsourcing platforms to annotate relevance labels for each query-document pair.
-Given the significant cost associated with human annotation, we adopt an alternative natural approach.
-Instead, we repurpose the original human-written corpus as our seed data source and directly instruct LLMs to rewrite each given text from the human-written corpus.
-As the written text generated by LLM carries almost the same semantic information as the original human-written text, we can assign the same relevance labels to new query-document pairs as those assigned to the original query-document pairs.
-Our instruction is straightforward: “Please rewrite the following text: {{human-written text}}”, as illustrated in the left part of Figure 2.
-This straightforward instruction enables LLMs to generate text without too many constraints while maintaining semantic equivalence to the original human-written text.
-Specifically, we choose Llama2 (Touvron et al., 2023) and ChatGPT to rewrite each seed human-written corpus, as Llama2 and ChatGPT are both the most widely-used and nearly the state-of-the-art open-sourced and closed-source LLM, respectively.
-We only generate texts with ChatGPT corresponding to the texts in SciFact dataset, mainly due to the significant cost involved in processing the larger NQ320K dataset.
-If not specified, the temperature of the used LLM for generation corpus is set as 0.2 by default for the following experiments.
-As a result, we can obtain two corresponding LLM-generated corpora with SciFact and NQ320K as seed data.
-After that, we extend the original labels of query and human-written text $\mathcal{R}^{H}=\{(q_{m},d^{H}_{m},r_{m})\}^{M}_{m=1}$ to get the corresponding label of LLM-generated text $\mathcal{R}^{G}=\{(q_{m},d^{G}_{m},r_{m})\}^{M}_{m=1}$.
-We will validate the quality of the datasets in the following section.
-Combining each original human-written corpus $\mathcal{C}^{H}$ with its corresponding LLM-generated corpus $\mathcal{C}^{G}$, original queries $\mathcal{Q}$, and labels $\mathcal{R}^{H}\bigcup\mathcal{R}^{G}$, we can create two new datasets, denoted as SciFact+AIGC and NQ320K+AIGC.
-Table 1 summarize the statistics of the proposed two datasets.
 […]
@@
 LLM-Generated 55.7 62.0 64.8 52.9 59.5 61.5 57.1 67.5 69.8 57.1 64.9 66.2
+2.2.1. Human-Written Corpus
+We first choose two widely used retrieval datasets written by humans in the Pre-LLM era as the seed data: SciFact and NQ320K.
+SciFact [^1] https://allenai.org/data/scifact (Wadden et al., 2020) dataset aims to retrieve evidence from the research literature containing scientific paper abstracts for fact-checking.
+NQ320K [^2] https://ai.google.com/research/NaturalQuestions (Kwiatkowski et al., 2019) is based on the Natural Questions (NQ) dataset from Google, where the documents are gathered from Wikipedia pages, and the queries are natural language questions.
+Following the practice in BEIR benchmark (Thakur et al., 2021), we process these two datasets in a standard format: corpus $\mathcal{C}^{H}$, queries $\mathcal{Q}$, and relevancy labels $\mathcal{R}^{H}=\{(q_{m},d^{H}_{m},r_{m})\}^{M}_{m=1}$, where $M$ is the number of labeled query-document pairs in the dataset.
+2.2.2. LLM-Generated Corpus
+For the LLM-generated corpus, we repurpose the original human-written corpus as our seed data and instruct LLMs to rewrite each given text from the human-written corpus.
+As the written text generated by LLM carries almost the same semantic information as the original human-written text, we can assign the same relevancy labels to new <query, LLM-generated document> pairs as those assigned to the original labeled <query, human-written document> pairs.
+Our instruction is straightforward: “Please rewrite the following text: {{human-written text}}”, as illustrated in the left part of Figure 2.
+This straightforward instruction enables LLMs to generate text without too many constraints while maintaining semantic equivalence to the original human-written text.
+Specifically, we choose Llama2 (Touvron et al., 2023) and ChatGPT to rewrite each seed human-written corpus, as Llama2 and ChatGPT are both the most widely-used and nearly the state-of-the-art open-sourced and closed-source LLM, respectively.
+We only generate texts with ChatGPT corresponding to the texts in SciFact dataset, mainly due to the significant cost involved in processing the larger NQ320K dataset.
+For the LLM-generated corpus, we conduct post-processing to remove unrelated parts of the original response from LLM like “Sure, here’s a possible rewrite of the text:”.
+As a result, we can obtain two corresponding LLM-generated corpora with SciFact and NQ320K as seed data.
+After that, we extend the original labels of query and human-written text $\mathcal{R}^{H}=\{(q_{m},d^{H}_{m},r_{m})\}^{M}_{m=1}$ to get the corresponding label of LLM-generated text $\mathcal{R}^{G}=\{(q_{m},d^{G}_{m},r_{m})\}^{M}_{m=1}$.
+We will validate the quality of the datasets in the following section.
+Combining each original human-written corpus $\mathcal{C}^{H}$ with its corresponding LLM-generated corpus $\mathcal{C}^{G}$, original queries $\mathcal{Q}$, and labels $\mathcal{R}^{H}\bigcup\mathcal{R}^{G}$, we can create two new datasets, denoted as SciFact+AIGC and NQ320K+AIGC.
+Table 1 summarizes the statistics of the proposed two datasets.
+(a) Jaccard Similarity
+(b) Overlap
+Figure 3. Distribution of term Jaccard similarity and overlap between Llama2-generated and human-written corpora.
 2.3. Statistics and Quality Validation of Datasets
 Take the Llama2-generated data as an example, we conduct the statistics and quality validation of the constructed datasets.
-The analysis of ChatGPT-generated datasets is provided in Appendix A.4, which shows similar observations and conclusions.
+The analysis of ChatGPT-generated datasets shows similar observations and conclusions and is omitted due to the page limitation.
 2.3.1. Term-based Statistics and Analysis
@@
 Specifically, we compute the Jaccard similarity ($\frac{|d^{G}\bigcap d^{H}|}{|d^{G}\bigcup d^{H}|}$) and the overlap ($\frac{|d^{G}\bigcap d^{H}|}{|d^{H}|}$) between each LLM-generated document and orginal human-written document.
-The distribution of these two term-based similarity metrics is shown in Figure 3.
-As we can see, both the Jaccard similarity and overlap distributions exhibit normal distribution, with peaks at about 0.4 and 0.6 for NQ320K+AIGC, and about 0.6 and 0.8 for SciFact+AIGC, respectively.
+As shown in Figure 3, both the Jaccard similarity and overlap distributions exhibit normal distribution, with peaks at about 0.6 and 0.8 for SciFact+AIGC, and about 0.4 and 0.6 for NQ320K+AIGC, respectively.
 These observations suggest that while there is a considerable overlap of terms between the LLM-generated text and the original human-written text, there are also distinct differences, especially noticeable in the NQ320K+AIGC dataset.
+(a) SciFact+AIGC
+(b) NQ320K+AIGC
+Figure 4. Semantic embedding visualization of different corpora on SciFact+AIGC and NQ320K+AIGC datasets.
 Figure 5. Distribution of cosine similarity of semantic embedding between Llama2-generated and human-written corpora.
@@
 For the LLM-generated texts, a pivotal consideration is whether they faithfully preserve the underlying semantics of the corresponding human-written corpus.
-If they indeed do so, we can confidently assign them the same relevance labels as the labels of their corresponding original human-written texts given each query.
-To assess this, we first leverage the OpenAI embedding model333text-embedding-ada-002:https://platform.openai.com/docs/api-reference/embeddings to acquire semantic embeddings for both the LLM-generated and human-written corpora.
+If they indeed do so, we then can confidently assign them the same relevancy labels as the labels of their corresponding original human-written texts given each query.
+To assess this, we first leverage the OpenAI embedding model333text-embedding-ada-002:https://platform.openai.com/docs/guides/embeddings to acquire semantic embeddings for both the LLM-generated and human-written texts.
 Subsequently, we visualize these embeddings through T-SNE (Van der Maaten and Hinton, 2008) in Figure 4.
-We observe a strikingly close overlap between the Llama2-generated corpus with the human-written corpus in the latent space.
-This observation strongly suggests that these Llama2-generated corpora adeptly preserve the original semantics.
+We observe a strikingly close overlap between the Llama2-generated corpus and the human-written corpus in the latent space.
+This observation strongly suggests that these LLM-generated corpora adeptly preserve the original semantics.
 Moreover, we delve into the cosine similarity of semantic embeddings between the LLM-generated text and their corresponding human-written counterparts.
 The results, as shown in Figure 5, also indicate a high degree of similarity, with most values exceeding 0.95, affirming the faithful preservation of semantics in LLM-generated text.
-Hence, for each query-document pair $(q,d^{G})$, we can confidently assign the relevant label $r$ to be the same as that of $(q,d^{H})$.
-To further validate the accuracy of the relevance label assignments, we conduct an evaluation of retrieval models on the human-written corpus and the LLM-generated corpus, respectively.
-The results are presented in Table 2.
-It is evident that all retrieval models exhibit no significant performance discrepancies in terms of various ranking metrics between the human-written and LLM-generated corpora across all datasets.
-This observation reinforces the confidence in the quality of our newly constructed datasets.
+Hence, for each query-document pair $(q,d^{G})$, we can confidently assign the relevancy label $r$ to be the same as that of $(q,d^{H})$.
+Table 3. Verification of semantics and text quality with human evaluation. The numbers in parentheses represent the proportion agreed upon by all three human annotators for each option.
+SciFact+AIGC NQ320K+AIGC
+Which document is more relevant to the given query?
+Human LLM Equal Human LLM Equal
+$0.0\%(0.0\%)$ $0.0\%(0.0\%)$ $100.0\%(82.0\%)$ $2.0\%(0.0\%)$ $0.0\%(0.0\%)$ $98.0\%(81.6\%)$
+Which document exhibits higher quality by considering the following aspects:
+linguistic fluency, logical coherence, and information density?
+Human LLM Equal Human LLM Equal
+$8.0\%(0.0\%)$ $6.0\%(0.0\%)$ $86.0\%(46.5\%)$ $4.0\%(0.0\%)$ $6.0\%(0.0\%)$ 90.0%(60.%)90.$0$ ( 60 . % )
 Table 4.
@@
 $\text{Relative}~{}~{}\Delta$25.5 -9.0 -5.5 -23.8 -10.7 -8.4 -22.6 -6.8 -6.5 -22.6 -9.3 -8.7
-3. RQ2: Explore The Source Bias
+2.3.3. Retrieval Performance Evaluation
+To further validate the accuracy of the relevancy label assignments, we conduct an evaluation of retrieval models on the human-written corpus and the LLM-generated corpus, respectively.
+The following representative retrieval models are adopted in the experiments: (1) Lexical Retrieval Models: TF-IDF (Sparck Jones, 1972) and BM25 (Robertson et al., 2009) and (2) Neural Retrieval Models: ANCE (Xiong et al., 2020), BERM (Xu et al., 2023), TAS-B (Hofstätter et al., 2021), Contriever (Izacard et al., 2021).
+The results on each sole source corpus on the proposed two new benchmarks are presented in Table 2.
+It is evident that all retrieval models exhibit no significant performance discrepancies in terms of various ranking metrics between the human-written and LLM-generated corpora across all datasets.
+This observation reinforces the confidence in the quality of our newly constructed datasets.
+2.3.4. Human Evaluation
+Note that in our constructed datasets, LLMs were instructed to rewrite human-written texts based solely on the original human-written text, without any query-related input, thereby preventing the additional query-specific information during rewriting.
+Moreover, to further verify this, we conduct a human evaluation.
+Specifically, we randomly select 50 <query, human-written document, LLM-generated document> triples from each dataset.
+The human annotators, comprising the authors and their highly educated colleagues, are asked to determine which document is more semantically relevant to the given query.
+The options are “Human”, “LLM”, or “Equal”.
+During the evaluation, annotators are unaware of the source of each document.
+Each triple is labeled at least by three different annotators, with the majority vote determining the final label.
+The results in Table 3, confirm that both sources of texts have almost the same semantic relevance to the given queries, which guarantees the fairness of our following exploration of source bias.
+Additionally, we also conduct further human evaluations specifically focused on text quality.
+The human annotators are asked to determine “Which document exhibits higher quality by considering the following aspects: linguistic fluency, logical coherence, and information density?” The notation process is the same as above, and the results are summarized in Table 3.
+The results indicate no significant distinction between LLM-generated and human-written content on text quality, demonstrating consistency across both sources.
+In fact, we also analyze the data cases and find that LLMs typically alter only parts of the vocabulary, leading to minor stylistic differences without impacting the core content, which can be further verified with these human evaluations.
+3. RQ2: Uncovering Source Bias
 In this section, we conduct extensive experiments on the constructed datasets to explore the source bias from various aspects.
@@
 We then conduct experiments with different retrieval models on both the first-stage retrieval and the second-stage re-ranking.
-Finally, we also conduct experiments on corpus generated by LLM with different temperatures to further investigate the source bias.
 3.1. Evaluation Metrics for Source Bias
-To quantitatively explore source bias, we calculate the ranking metrics, targeting separately either human-written or LLM-generated corpus.
-Specifically, for each query, an IR model produces a ranking list that comprises documents from both human-written and LLM-generated corpora.
-We then calculate ranking metrics separately for human-written and LLM-generated texts, depending on the target data source.
-When computing metrics for one target (e.g., human-written corpus $\mathcal{C}^{H}$), the data corresponding to the other side (e.g., LLM-generated corpus $\mathcal{C}^{G}$) is masked or disregarded (i.e., mask all the positive label $r\in\mathcal{R}^{G}$ as negative), but the rank of each document is still based on the original ranking list that incorporates a mixture of both types of text.
-This approach allows us to independently assess the performance of IR models on different corpus sources.
-For the ranking metrics, we report the top-$K$ Normalized Discounted Cumulative Gain (NDCG@$K$) and Mean Average Precision (MAP@$K$) with $K\in\{1,3,5\}$.
-For a better understanding of the calculation of the evaluation metrics, we provide a toy example in Appendix A.1.
+To quantitatively explore source bias, we calculate ranking metrics, targeting separately either human-written or LLM-generated corpus.
+Specifically, for each query, an IR model produces a ranking list that comprises documents from mixed corpora.
+We then calculate top-$K$ Normalized Discounted Cumulative Gain (NDCG@$K$) and Mean Average Precision (MAP@$K$), for $K\in\{1,3,5\}$, independently for each corpus source.
+When assessing one corpus (e.g., human-written), documents from the other (e.g., LLM-generated) are treated as non-relevant, though the original mixed-source ranking order is maintained.
+This approach allows us to independently assess the performance of IR models on each corpus source.
 To better normalize the difference among different benchmarks, we also introduce the relative percentage difference as follows:
@@
 Note that Relative $\Delta>0$ means retrieval models rank human-written texts higher, and Relative $\Delta<0$ indicates LLM-generated texts are ranked higher.
-The greater the absolute value of Relative $\Delta$, the greater the ranking performance difference on two corpora.
-3.2. Bias in Different Retrieval Models
-Our initial set of experiments is aimed at conducting a comprehensive analysis of different retrieval models within the context of the LLM era.
-These models encompass both traditional lexical approaches and modern neural IR models.
-The results targeted on either human-written or LLM-generated corpus on the SciFact+AIGC and NQ320K+AIGC are reported in Table 3.
-Based on these results, we have the following noteworthy observations and conclusions:
-Lexical approaches tend to favor human-written documents.
-When the retrieved corpora contain both human-written and LLM-generated texts, TF-IDF and BM25 achieve better results on the human-written corpus.
-This phenomenon is observed in terms of almost all ranking metrics across both the SciFact+AIGC and NQ320K+AIGC datasets.
-A plausible explanation for this phenomenon lies in the term-based distinctions between text generated by LLMs and human-written content, as evident in Figure 3.
-On the other hand, the queries are crafted by humans and thus exhibit a style more closely aligned with human-written text.
-Neural retrieval models are biased towards LLM-generated texts.
-An intriguing phenomenon emerges when evaluating neural retrieval models that rely on semantic matching.
-There is a significant performance gap between their performance on the human-written corpus and LLM-generated corpus.
-Note that all neural retrieval models achieve a relative performance gain of over $30\%$ when targeting LLM-generated corpus as compared to human-written corpus.
-These findings suggest an inherent bias in neural retrieval models towards LLM-generated text, which we named the source bias.
-This bias may arise from the potential exploitation of specific patterns or ‘semantic shortcuts’ embedded in LLM-generated text, given that both neural retrieval models (based on PLMs) and LLMs share a similar Transformer-based (Vaswani et al., 2017) architecture and a similar pretraining paradigm.
-Moreover, it seems that the information in texts is compressed at the semantic level during LLM generation, resulting in a clearer understanding by neural retrieval models.
-We will delve deeper into the cause of this bias in the subsequent section.
-This set of experiments provides valuable insights into the performance of different retrieval models in the LLM era, underscoring the presence of the source bias within neural retrieval models.
-Next, we will further uncover more scenarios that exhibit this bias.
-Table 4.
-Performance comparison of different neural retrieval models for mixed human-written and ChatGPT-generated corpora on SciFact+AIGC dataset.
-The annotations in the table are consistent with Table 3.
+The greater the absolute value of Relative $\Delta$, the greater the ranking performance difference between two sourced content.
+3.2. Bias in Neural Retrieval Models
+Table 5. Performance comparison of different neural retrieval models for mixed human-written and ChatGPT-generated corpora on SciFact+AIGC dataset.
 Model Target Corpus NDCG@1 NDCG@3 NDCG@5 MAP@1 MAP@3 MAP@5
@@
 $\text{Relative}~{}~{}\Delta$31.6 -9.3 -7.0 -29.3 -11.8 -10.2
-Table 5.
-Bias evaluation of re-ranking models on mixed human-written and Llama2-generated corpora of SciFact+AIGC dataset.
-The re-ranking methods rerank the top-100 retrieved hits from a first-stage BM25 model.
-The annotations in the table are consistent with Table 3.
-Metrics Target Corpus Base Re-ranking
-BM25 +MiniLM +monoT5
-NDCG@1 Human-Written 26.7 21.3 19.7
-LLM-Generated 21.0 32.7 39.7
-$\text{Relative}~{}~{}\Delta$ 23.9 -42.2 -67.3
-NDCG@3 Human-Written 40.3 42.8 45.9
-LLM-Generated 38.8 47.8 52.9
-$\text{Relative}~{}~{}\Delta$ 3.8 -11.0 -14.2
-NDCG@5 Human-Written 44.4 46.9 49.0
-LLM-Generated 41.5 50.2 54.7
-$\text{Relative}~{}~{}\Delta$ 6.8 -6.8 -11.0
-MAP@1 Human-Written 25.7 20.8 18.9
-LLM-Generated 19.6 30.8 37.8
-$\text{Relative}~{}~{}\Delta$ 26.9 -38.8 -66.7
-MAP@3 Human-Written 36.7 37.5 39.7
-LLM-Generated 34.3 43.6 48.9
-$\text{Relative}~{}~{}\Delta$ 6.8 -15.0 -20.8
-MAP@5 Human-Written 39.1 40.0 41.6
-LLM-Generated 35.9 45.0 50.1
-$\text{Relative}~{}~{}\Delta$ 8.5 -11.8 -18.5
-3.3. Bias in ChatGPT-Generated Corpus
-To strengthen our argument that the source bias is not limited to any specific LLM-generated corpus, we extend our investigation to include ChatGPT, another widely adopted and nearly state-of-the-art LLM.
+In our assessment of various retrieval models on SciFact+AIGC and NQ320K+AIGC datasets, we observe distinct phenomena when evaluating against human-written and LLM-generated corpora, as reported in Table 4.
+Our key findings are as follows:
+Lexical models prefer human-written texts.
+Lexical models like TF-IDF and BM25 show a tendency to favor human-written texts over LLM-generated texts across most ranking metrics in both datasets.
+A plausible explanation for this phenomenon lies in the term-based distinctions between text generated by LLMs and human-written content, as evident in Figure 3.
+Additionally, the queries are crafted by humans and thus exhibit a style more closely aligned with human-written text.
+Neural retrievers are biased towards LLM-generated texts.
+Neural models, which rely on semantic matching with PLMs, demonstrate a pronounced preference for LLM-generated texts, often performing over 30% better on these compared to human-written texts.
+These findings suggest an inherent bias in neural retrievers towards LLM-generated text, which we named the source bias.
+This source bias may stem from PLMs-based neural retrievers and LLMs sharing similar Transformer-based architectures (Vaswani et al., 2017) and pretraining approaches, leading to potential exploitation of semantic shortcuts in LLM-generated text during semantic matching.
+Additionally, LLMs seem to semantically compress information in a manner that makes it more comprehensible to neural models.
+A deeper exploration into the causes of source bias is presented in the following section.
+To strengthen our conclusion that source bias is not limited to any specific LLM, we extend our investigation to include ChatGPT, another widely adopted and nearly state-of-the-art LLM.
 We employ ChatGPT to generate a corpus using the same prompts as those utilized with Llama2 in the above experiments.
-In Appendix A.4, we present the detailed data quality evaluation of the ChatGPT-generated corpus.
 Subsequently, in Table 5, we report the evaluation results on the SciFact+AIGC dataset, which contains both human-written and ChatGPT-generated texts.
@@
 This observation provides additional substantiation of the presence of source bias within these neural retrieval models.
-Table 6.
-Bias evaluation of re-ranking models on mixed human-written and ChatGPT-generated corpora of SciFact+AIGC dataset.
-The re-ranking methods rerank the top-100 retrieved hits from a first-stage BM25 model.
-The annotations in the table are consistent with Table 3.
-Metrics Target Corpus Base Re-ranking
-BM25 +MiniLM +monoT5
-NDCG@1 Human-Written 24.3 18.3 21.3
-LLM-Generated 24.3 35.7 39.3
-$\text{Relative}~{}~{}\Delta$ 0.0 -64.4 -59.4
-NDCG@3 Human-Written 38.5 41.4 46.4
-LLM-Generated 40.2 50.1 54.2
-$\text{Relative}~{}~{}\Delta$4.3 -19.0 -15.5
-NDCG@5 Human-Written 42.7 45.6 48.9
-LLM-Generated 42.7 53.0 56.1
-$\text{Relative}~{}~{}\Delta$ 0.0 -15.0 -13.7
-MAP@1 Human-Written 23.7 17.9 20.5
-LLM-Generated 23.1 33.8 37.8
-$\text{Relative}~{}~{}\Delta$ 2.6 -61.5 -59.3
-MAP@3 Human-Written 34.8 35.8 40.3
-LLM-Generated 35.8 45.9 50.0
-$\text{Relative}~{}~{}\Delta$2.8 -24.7 -21.5
-MAP@5 Human-Written 37.3 38.3 41.7
-LLM-Generated 37.3 47.6 51.4
-$\text{Relative}~{}~{}\Delta$ 0.0 -21.7 -20.8
+Furthermore, we also explore the popular InstructGPT-prompts GitHub Repository, which includes several common prompts for rephrasing passages  [^4] https://github.com/kevinamiri/Instructgpt-prompts.
+The experimental results in Appendix B show varying degrees of source bias, indicating that common prompts can easily trigger source bias with LLM-generated content.
+These findings highlight the notable presence of source bias in neural retrieval models towards LLM-generated content.
+Table 6. Bias evaluation of re-ranking models on SciFact+AIGC dataset. The re-ranking methods rerank the top-100 retrieved hits from a first-stage BM25 model.
+Metrics Target Corpus Llama2-generated ChatGPT-generated
+BM25 +MiniLM +monoT5 BM25 +MiniLM +monoT5
+NDCG@1 Human-Written 26.7 21.3 19.7 24.3 18.3 21.3
+LLM-Generated 21.0 32.7 39.7 24.3 35.7 39.3
+$\text{Relative}~{}~{}\Delta$ 23.9 -42.2 -67.3 0.0 -64.4 -59.4
+NDCG@3 Human-Written 40.3 42.8 45.9 38.5 41.4 46.4
+LLM-Generated 38.8 47.8 52.9 40.2 50.1 54.2
+$\text{Relative}~{}~{}\Delta$ 3.8 -11.0 -14.2 -4.3 -19.0 -15.5
+NDCG@5 Human-Written 44.4 46.9 49.0 42.7 45.6 48.9
+LLM-Generated 41.5 50.2 54.7 42.7 53.0 56.1
+$\text{Relative}~{}~{}\Delta$ 6.8 -6.8 -11.0 0.0 -15.0 -13.7
+MAP@1 Human-Written 25.7 20.8 18.9 23.7 17.9 20.5
+LLM-Generated 19.6 30.8 37.8 23.1 33.8 37.8
+$\text{Relative}~{}~{}\Delta$ 26.9 -38.8 -66.7 2.6 -61.5 -59.3
+MAP@3 Human-Written 36.7 37.5 39.7 34.8 35.8 40.3
+LLM-Generated 34.3 43.6 48.9 35.8 45.9 50.0
+$\text{Relative}~{}~{}\Delta$ 6.8 -15.0 -20.8 -2.8 -24.7 -21.5
+MAP@5 Human-Written 39.1 40.0 41.6 37.3 38.3 41.7
+LLM-Generated 35.9 45.0 50.1 37.3 47.6 51.4
+$\text{Relative}~{}~{}\Delta$ 8.5 -11.8 -18.5 0.0 -21.7 -20.8
 3.3. Bias in Re-Ranking Stage
@@
 The first stage involves a retriever, responsible for document recall, while the second stage employs a re-ranker, which fine-tunes the ordering of documents within the initially retrieved set.
-Our investigation has revealed the presence of the source bias in the initial document filtering process (i.e., retrieval stage).
-Hence, a natural pivotal research question arises: does this bias extend to the ranking models used in the second-stage document filtering process (re-ranking stage), within IR systems?
-To delve into this, we choose two representative re-ranking models: MiniLM (Wang et al., 2020) and monoT5 (Nogueira et al., 2020) to rerank the top-100 document list retrieved by a first-stage BM25 model.
-MiniLM is a cross-encoder model pre-trained on MS MARCO (Nguyen et al., 2016) using a knowledge distillation on three teacher models: BERT-base (Devlin et al., 2019), BERT-large (Devlin et al., 2019), and ALBERT-large (Lan et al., 2019) models.
-monoT5 is a sequence-to-sequence re-ranker that uses T5 (Raffel et al., 2020) to calculate the relevance score, which is also pretrained on MS MARCO.
-The results on the SciFact+AIGC dataset with Llama-generated corpus and ChatGPT-generated corpus are presented in Table 5 and Table 6, respectively.
+While we have revealed the presence of the source bias in the first stage, a natural pivotal research question remains: does this bias also manifest in the re-ranking stage?
+To delve into this, we select two representative and state-or-the-art re-ranking models: MiniLM (Wang et al., 2020) and monoT5 (Nogueira et al., 2020) to rerank the top-100 document list retrieved by a first-stage BM25 model.
+The results on the SciFact+AIGC dataset with Llama-generated corpus and ChatGPT-generated corpus are presented in Table 6.
 From the results, while even the first-stage retrievers (BM25) may exhibit a preference for human-written content, the second-stage re-rankers once again demonstrate a bias in favor of LLM-generated content.
-Remarkably, the bias in re-ranking models appears to be more pronounced, as evidenced by the relative percentage difference of $67.3\%$ and $59.4\%$ in NDCG@$1$ for monoT5 in Table 5 and Table 6, respectively.
-These findings further solidify the pervasiveness of source bias in neural ranking models that rely on PLMs, regardless of the retrieval stage or re-ranking stage.
-(a) Relative $\Delta$ on NDCG@$1$
-(b) Relative $\Delta$ on MAP@$1$
-Figure 6.
-Performance comparison (NDCG@$1$ and MAP@$1$) of neural retrieval models on SciFact+AIGC dataset under different temperatures of Llama2-generated corpus.
-A larger negative the Relative $\Delta$ indicates a more severe source bias.
-3.5. Bias in Different Temperatures
-Temperature is an important hyperparameter used to control the randomness and creativity of text generation in LLMs.
-A lower temperature sharpens the probability distribution of predicted words, favoring the most likely words.
-Conversely, a higher temperature flattens the probability distribution, allowing all words to have a more equal chance of being selected.
-To investigate the impact of temperature on the source bias, we conduct experiments in which Llama2 generated different corpora with varying temperatures (i.e., $T\in\{0.2,1.0\}$).
-We then examine different neural retrieval models by combining the Llama2-generated corpus with the human-written corpus one by one.
-The relative gaps in NDCG@$1$ and MAP@$1$ on the SciFact+AIGC dataset are presented in Figure 6.
-In Appendix A.4, we present detailed data statistics and quality evaluation of the Llama2-generated corpus ($T=1.0$).
-We also provide the raw results of these retrieval models for the mixed human-written and Llama2-generated corpora ($T=1.0$) in  Table 9 of Appendix  A.5.
-The results suggest that all neural retrieval models still exhibit a bias towards LLM-generated text, regardless of the temperature.
-Furthermore, there is a reduced bias on most neural retrieval models when applied to LLM-generated text with higher temperatures, especially on the re-ranking models.
-Note that if the temperature coefficient is higher ($T>1.0$), the LLM will not be able to generate coherent meaningful texts, and the generated results will be very random.
-These findings provide us with potential starting points for analyzing this bias and we will illustrate this in the next section.
+Remarkably, the bias in re-ranking models appears to be more severe, as evidenced by the relative percentage difference of $-67.3\%$ and $-59.4\%$ in NDCG@$1$ for monoT5, respectively.
+These findings further confirm the pervasiveness of source bias in neural ranking models that rely on PLMs, regardless of the first retrieval stage or second re-ranking stage.
 4. RQ3: The Cause of Source Bias
-In this section, we delve deeper into the cause of the observed source bias.
-We aim to confirm whether LLM-generated texts, having less noise and more concentrated semantic topics, become easier for neural retrieval models to comprehend.
-We conduct a series of analyses from the perspective of text compression and provide valuable theoretical insights.
+In this section, we delve deeper into why neural retrieval models exhibit source bias.
+Our objective is to determine whether the LLM-generated texts, characterized by reduced noise and more concentrated semantic topics, are inherently easier for neural retrieval models to semantically match.
+We conduct a series of analyses from the perspective of text compression and provide valuable insights.
 Figure 6. Comparision of the relative singular value (SV) of the different corpus after SVD. The singular values are sorted in descending order from left to right.
-Figure 8. Comparision of the relative singular value of LLM-generated corpus with different temperatures. The singular values are sorted in descending order from left to right.
-4.1. Viewpoint from Compression
-In this analysis, we explore the cause of source bias through a compression perspective, drawing inspiration from recent studies that suggest LLMs act as lossless compressors (Delétang et al., 2023).
-We postulate that LLMs consistently focus on the major part of the information while filtering out the noise.
-This, however, does not hold true for humans, as human minds tend to get distracted by chance occurrences.
-To conduct this examination, we assess the number and strength of topics in both the human-written corpus and the LLM-generated corpus.
-We employ Singular Value Decomposition (SVD) (Klema and Laub, 1980) on these two corpora.
+4.1. Viewpoint from Text Compression
+We first explore the cause of source bias from a compression perspective, drawing inspiration from recent studies that suggest LLMs are lossless compressors (Delétang et al., 2023).
+We hypothesize that LLMs efficiently focus on essential information, minimizing noise during generation, in contrast to human-written texts, which may include more diverse topics and incidental noise.
+To verify this, we employ Singular Value Decomposition (SVD) (Klema and Laub, 1980) to compare topic concentration and noise in human-written and LLM-generated texts.
 The dimension of the SVD corresponds to the maximum number of topics, and the singular value associated with each topic represents its strength.
 High singular values predominantly capture primary topic information, whereas low singular values indicate noise.
-To be more specific, we utilize the OpenAI embedding model to obtain embedding matrices for each corpus within the SciFact+AIGC dataset.
-Subsequently, we perform SVD on these matrices.
-The singular values are arranged in descending order, and their comparison to the human-written corpus is visualized in Figure 7.
-As illustrated in Figure 7, LLM-generated texts exhibit larger singular values at the top large singular values, while smaller singular values at the tail small singular values.
-This observation suggests that LLM-generated texts tend to have more focused semantics with less noise, rendering them more suitable for precise semantic similarity calculations.
-In contrast, human-written texts often contain a wider range of topics and higher levels of noise, making them harder for neural IR models to understand.
-As a result, this difference in semantic concentration may contribute to the observed bias in neural IR models.
-Figure 9. Comparision of the PPL of the different corpora.
-Furthermore, we conduct an SVD analysis on LLM-generated documents with different temperature coefficients, as shown in Figure 8.
-Evidently, for various temperature coefficients of Llama2, the head of the SVD consistently displays higher singular values compared to human-written text.
-However, the tails of the SVD corresponding to higher temperature coefficients align more closely with the human-written text, while lower temperature coefficients result in tails with smaller singular values.
-This aligns with our previous findings in Section 3.5, that is, lower temperature results in more conservative and predictable text, and higher temperature results in more creative and diverse text.
-Thus, lower temperature coefficients lead to more substantial compression, greater topic concentration, and a more pronounced source bias.
+Specifically, we utilize the OpenAI embedding model to obtain embedding matrices for each corpus in the SciFact+AIGC dataset and then conduct SVD.
+The resulting singular values are arranged in descending order, and their comparison to the human-written corpus is visualized in Figure 6.
+As we can see, LLM-generated texts exhibit larger singular values at the top large singular values, while smaller singular values at the tail small singular values.
+This observation suggests that LLM-generated texts tend to have more focused semantics with less noise, rendering them more suitable for precise semantic matching.
+In contrast, human-written texts often contain a wider range of latent topics and higher levels of noise, making them harder for neural retrievers to understand.
+As a result, this difference in semantic concentration may contribute to the observed source bias in neural retrievers.
+Figure 7. Comparision of the PPL of the different corpus.
 4.2. Further Analysis from Perplexity
-Considering that most modern neural IR models are grounded on PLMs (Yates et al., 2021; Guo et al., 2020; Zhao et al., 2022), such as BERT (Devlin et al., 2019), Roberta (Liu et al., 2019), and T5 (Raffel et al., 2020), we try to analyze the perplexity of PLMs to further support the conclusion above from the viewpoint of compression that LLM-generated texts can be better understood by PLMs.
+Considering that most modern neural retrievers are grounded on PLMs (Yates et al., 2021; Guo et al., 2020; Zhao et al., 2023a), such as BERT (Devlin et al., 2019), Roberta (Liu et al., 2019), and T5 (Raffel et al., 2020), we analyze the perplexity of PLMs to further support the conclusion above from the viewpoint of compression that LLM-generated texts can be better understood by PLMs.
 Perplexity is an important metric for evaluating how well a language model can understand a given text (Azzopardi et al., 2003; Wang et al., 2019).
@@
 $\text{PPL}(d)=-\frac{1}{S}\left(\sum_{s=1}^{S}\log P_{\text{LM}}(d_{s}|\text{context})\right),$
-where $S$ represents the token length of each text $d$ and $P_{\text{LM}}(d_{s})$ is the predicted likelihood of the $s$-th token conditioned on the context.
-Lower perplexity suggests more confidence and understanding of LM to text patterns, while higher perplexity implies greater uncertainty in predictions, often arising from complex or unpredictable text patterns.
-Using the most widely-used LM, BERT (Devlin et al., 2019), as an example, we employ it to calculate the PPL for various corpora.
-As BERT is not an autoregressive LM, we follow standard practices (Wang et al., 2021; Wang and Cho, 2019) to calculate the predicted likelihood of each token conditioned on the other tokens, i.e.,
+where $S$ is the token length of text $d$ and $P_{\text{LM}}(d_{s})$ is the predicted likelihood of the $s$-th token conditioned on the context.
+Lower perplexity suggests more confidence and understanding of LM for text patterns, while higher perplexity implies greater uncertainty in predictions, often arising from complex or unpredictable text patterns.
+Using the most widely-used LM, BERT (Devlin et al., 2019), as an example, we employ it to calculate the PPL for different corpus.
+As BERT is not an autoregressive LM, we follow standard practices (Wang et al., 2021; Wang and Cho, 2019) to calculate the likelihood of each token conditioned on the other tokens, i.e.,
 $P_{\text{LM}}(d_{s}|\text{context}):=P_{\text{BERT}}(d_{s}|d_{\leq S\backslash\{s\}}).$
-The distribution of perplexity for different corpus within the SciFact+AIGC dataset is shown in Figure 9.
-Notably, LLM-generated texts consistently exhibit significantly lower perplexity (Figure 9), implying enhanced comprehensibility and higher confidence from BERT’s perspective.
-Consequently, neural retrieval models based on PLMs can more effectively comprehend LLM-generated texts.
-This enhanced understanding allows for a more accurate prediction of relevance, leading to the observed source bias in favor of LLM-generated texts.
-Moreover, as shown in Figure 9, Llama2 with a lower temperature coefficient (i.e., $T=0.2$) generates text with a lower perplexity than the higher temperature coefficient (i.e., $T=1.0$).
-This observation indicates that LLM with a smaller temperature is less prone to generate unexpected or unusual words, thereby enabling a more effective understanding and grasp of language patterns by the PLMs.
-To conclude, these observations underscore that texts generated by LLM are more easily comprehensible when processed by PLMs in neural IR models, culminating in the observed source bias.
-4.3. Theoretical Analysis and Insights
-In  Figure 9, we have compared the PPL for different corpus using the BERT model.
-In this section, we aim to further provide some theoretical insights into the above observations that LLM-generated texts have a smaller perplexity than human-written texts.
-Without loss of generality, we define the PPL in an autoregressive manner.
-Let $d^{H}$ denote a document written by humans, and $d^{G}$ a document generated by an LLM conditioned on $d^{H}.$ For a given document $d$ and BERT model $\mathcal{B}$, PPL is calculated as
-$\text{PPL}(d,\mathcal{B})=-\frac{1}{S}\sum_{s=1}^{S}\log P_{\text{BERT}}(d_{s}|d_{<{s}}).$
-Similarly, we use $\text{PPL}(d,\mathcal{H})$ to represent the PPL of document $d$ when evaluated by humans. The PPL of $d^{G}$ conditioned on $d^{H}$ is denoted as
-$\text{PPL}(d^{G}\mid d^{H},\mathcal{B})=-\frac{1}{S}\sum_{s=1}^{S}\log P_{\text{BERT}}(d_{s}^{G}|d_{<{s}}^{G},d^{H}).$
-When evaluated by humans, we use $\text{PPL}(d^{G}\mid d^{H},\mathcal{H})$ to represent the PPL of $d^{G}$ conditioned on $d^{H}$.
-In the theorem below, we introduce two assumptions: Semantic Superiority and Conditional Redundancy, to theoretically establish the sufficient conditions under which $\text{PPL}(d^{G},\mathcal{B})\leq\text{PPL}(d^{H},\mathcal{B})$ holds.
-Semantic Superiority suggests that the perplexity of human-written texts, when evaluated by humans, is lower than when evaluated by BERT.
-Conditional Redundancy implies that the perplexity of $d^{G}$, given $d^{H}$, is lower than the perplexity of $d^{H}$ when evaluated directly.
-This is intuitively true when the information added in generating from $d^{H}$ to $d^{G}$ doesn’t exceed the original information in $d^{H}$.
-Then we have the following theorem:
-Theorem 4.1 ().
-Given the following conditions:
-  • [leftmargin=0.5cm]
-  • Semantic Superiority: human beings outperform BERT in understanding human-written texts, i.e.,
-    $\textnormal{PPL}(d^{H},\mathcal{B})-\textnormal{PPL}(d^{H},\mathcal{H})\geq 0.$
-  • Conditional Redundancy: generating $d^{G}$ from $d^{H}$ adds less perplexity than $d^{H}$ itself, i.e.,
-    $\textnormal{PPL}(d^{H},\mathcal{H})-\textnormal{PPL}(d^{G}\mid d^{H},\mathcal{H})\geq 0.$
-  • Bounded Perplexity: there exists a bounded non-negative difference $\epsilon$ in BERT’s perplexity for $d^{G}$ with or without $d^{H},$ i.e.,
-    $\textnormal{PPL}(d^{G},\mathcal{B})-\textnormal{PPL}(d^{G}\mid d^{H},\mathcal{B})\leq\epsilon.$
-if LLM aligns more closely with BERT than with human when predicting $d^{G}$ given $d^{H}$, such that for any $s\in[S],$
-[…]
-[…]
-it follows that
-[…]
-Proof of Theorem 4.1 can be found in Appendix A.3.
-In  Theorem 4.1, we assume that there exists an upper bound $\epsilon$ on the increase in perplexity when evaluating $d^{G}$directly, compared to evaluating $d^{G}$ conditioned on $d^{H}$.
-Moreover, the KL divergence is used to compare the distributions of the document $d^{G}$ conditioned on $d^{H}$ according to the LLM, BERT model, and humans.
-It is worth to emphasize that inequation  (1) is not the assumption on the understanding capabilities of BERT, LLM, and humans.
-Instead, this inequation assumes that when predicting $d^{G}$ given $d^{H},$ the predictions by LLM are more closely aligned with those of BERT.
-We demonstrate that, when inequation  (1) is satisfied, the perplexity(evaluated by PLMs such as BERT) of $d^{G}$ is lower than that of $d^{H}$.
-We’d like to emphasize that it is reasonable to expect that inequation  (1) holds true because both LLM and BERT are Transformer-based models that use similar pretraining paradigms.
-The commonality in model structure and learning paradigms may lead to similar inherent biases in text prediction, making their predictions more aligned with each other.
+The distribution of perplexity for different corpus in the SciFact+AIGC dataset is shown in Figure 7.
+Notably, LLM-generated texts consistently exhibit significantly lower perplexity, indicating enhanced comprehensibility and higher confidence from BERT’s perspective.
+Consequently, PLMs-based neural retrievers can more effectively model the semantics of LLM-generated texts, leading to the observed source bias in favor of LLM-generated texts.
+In Appendix C, we also provide a theoretical analysis to illustrate and verify the observation in Figure 7 that LLM-generated texts have a smaller perplexity than human-written texts.
+5. RQ4: Mitigating Source Bias
+In this section, we propose a simple but effective approach to mitigate source bias by introducing a debiased constraint to the optimization objective.
+In this way, we can force the neural IR models to focus on modeling semantic relevance rather than the inherent semantic shortcut of the LLM-generated content.
 (a) Results on mixed human-written and Llama2-generated corpora
 (b) Results on mixed human-written and ChatGPT-generated corpora
-Figure 10.
-Performance comparison (NDCG@$1$ and MAP@$1$) of neural retrieval models on SciFact+AIGC dataset with different debiased co-efficient $\alpha$ in our proposed debiasing method.
-A larger positive the Relative $\Delta$ indicates retrieval models rank human-written text more higher than LLM-generated text.
-The grey dashed line represents Relative $\Delta=0$.
-(a) ANCE
-(b) BERM
-Figure 11.
-Performance comparison of neural retrievers on only human-written SciFact dataset with different co-efficient $\alpha$ in our proposed debiased method.
-The results on MAP have a similar tendency and are omitted for simplicity.
-5. RQ4: Mitigating Source Bias
-Our previous extensive experiments and analysis have revealed that the source bias widely exists in different neural retrievers due to that neural retrievers appear to adopt inherent semantic shortcuts when learning the semantic representation of LLM-generated content.
-In this section, we propose a simple but effective approach to mitigate source bias by introducing a debiased constraint to the optimization objective.
-In this way, we can force the neural IR models to focus on modeling semantic relevance rather than the inherent semantic shortcut of the LLM-generated content.
+Figure 8. Performance comparison (NDCG@$1$) of neural models on SciFact+AIGC dataset with different debiased co-efficient $\alpha$. The grey dashed line represents Relative $\Delta=0$.
 5.1. Our Method: A Debiased Constraint
-Our earlier findings of source bias indicate that neural retrieval models tend to rank LLM-generated documents in higher positions.
-Thus, the motivation of our debiased method is straightforward, which is to force the retrieval models to focus on modeling the semantic relevance and not to assign higher predicted relevance scores to the LLM-generated documents.
+Our earlier findings of source bias indicate that neural retrievers tend to rank LLM-generated documents in higher positions.
+Thus, the motivation of our debiased method is straightforward, which is to force the retrieval models to focus on modeling the semantic relevance and not assign higher predicted relevance scores to the LLM-generated documents.
 Specifically, following the practice in Section 2.2, we first generate the corresponding LLM-generated corpus $\mathcal{C}^{G}$ for the original human-written training corpus $\mathcal{C}^{H}$.
-In this way, we can get the new paired training data $\mathcal{D}=\{(q_{m},d^{H}_{m},d^{G}_{m})\}_{m=1}^{M}$, where each element $(q_{m},d^{H}_{m},d^{G}_{m})$ is a query, human-written document, and LLM-generated document triplet.
+In this way, we can get the new paired training data $\mathcal{D}=\{(q_{m},d^{H}_{m},d^{G}_{m})\}_{m=1}^{M}$, where each element $(q_{m},d^{H}_{m},d^{G}_{m})$ is a <query, human-written document, LLM-generated document> triplet.
 $d^{H}_{m}$ and $d^{G}_{m}$ are the corresponding human-written and LLM-generated relevant documents for the query $q$, respectively.
@@
 (1) $\mathcal{L}_{\text{debias}}=\sum_{(q_{m},d^{H}_{m},d^{G}_{m})\in\mathcal{D}}\max\{0,\hat{r}(q,d^{G};\Theta)-\hat{r}(q,d^{H};\Theta)\}$
-where the $\hat{r}(q,d^{G};\Theta)$ and $\hat{r}(q,d^{H};\Theta)$ are the predicted relevance score of $(q,d^{G})$ and $(q,d^{H})$ by the retrieval models with parameters $\Theta$, respectively.
-This constraint can measure and penalize the source bias when the predicted relevance score of $(q,d^{G})$ is greater than that of $(q,d^{H})$.
-Therefore, we can mitigate the source bias by minimizing this debiased constraint.
-Based on the debiased constraint defined in (2), we can define the final loss for training an unbiased neural retrieval model:
+where $\hat{r}(q,d^{G};\Theta)$ and $\hat{r}(q,d^{H};\Theta)$ are the predicted relevance scores of $(q,d^{G})$ and $(q,d^{H})$ by the retrieval models with parameters $\Theta$, respectively.
+This constraint can penalize biased samples when the predicted relevance score of $(q,d^{G})$ is greater than that of $(q,d^{H})$.
+Based on the debiased constraint defined in (1), we can define the final loss for training an unbiased neural retriever:
 (2) $\mathcal{L}=\mathcal{L}_{\text{rank}}+\alpha\mathcal{L}_{\text{debias}}$
-where the $\mathcal{L}_{\text{rank}}$ can be any common-used loss for the ranking task, e.g., contrastive loss or regression loss (Zhao et al., 2022; Guo et al., 2020, 2022).
+where the $\mathcal{L}_{\text{rank}}$ can be any common-used loss for the ranking task, e.g., contrastive loss or regression loss (Zhao et al., 2023a; Guo et al., 2020, 2022).
 And $\alpha$ is the debiased co-efficient that can balance the ranking performance and the degree of the source bias.
-The larger $\alpha$ indicates the greater penalty on the samples with source bias, leading to the retriever being more likely to rank the human-written text to a higher position.
+The larger $\alpha$ indicates the greater penalty on the biased samples, leading to the retriever being more likely to rank the human-written texts in higher positions.
 5.2. Results and Analysis
 To evaluate the effectiveness of our proposed debiased method, we equip the debiased constraint defined in Eq.
-(2) to two representative neural retrieval models, ANCE (Xiong et al., 2020) and BERM (Xu et al., 2023).
+(1) to two representative neural retrievers: ANCE (Xiong et al., 2020) and BERM (Xu et al., 2023).
 In the experiments, we vary the debiased co-efficient $\alpha$ within the range of $\{1e\text{-}4,5e\text{-}4,1e\text{-}3,5e\text{-}3,1e\text{-}2\}$.
-The original retrieval models learned without our proposed debiased constraint are denoted as “w/o debias”.
-The results for both mixed human-written and Llama2-generated corpora, and mixed human-written and ChatGPT-generated corpora of SciFact+AIGC dataset are presented in Figure 10.
+The original retrieval models learned without the debiased constraint are denoted as “w/o debias”.
+The results on the SciFact+AIGC dataset are presented in Figure 8.
+(a) ANCE
+(b) BERM
+Figure 9. Performance comparison of neural retrievers on only human-written SciFact dataset with different co-efficient $\alpha$ in our proposed debiased method.
 As we can see, as the debiased co-efficient $\alpha$ increases, the Relative $\Delta$ gradually shifts from negative to positive across almost all metrics and mixed datasets.
@@
 Moreover, as shown in Figure 9, our method not only maintains the retrieval performance on the sole human-written corpus but also provides improvements, especially with BERM as the backbone.
-The reason can be that the introduction of many LLM-generated samples by our method may enhance the retrieval model’s ability to estimate relevance for distinguishing similar samples.
-In summary, these empirical results have demonstrated the efficacy of our proposed debiased method in mitigating the source bias to different degrees through the adjustment of the debiased coefficient $\alpha$.
-This flexibility enables us to customize our debiasing mechanism to meet diverse perspectives and demands.
+This improvement is likely due to the inclusion of LLM-generated samples, which might enhance the model’s ability to discern relevance among similar documents.
+In summary, these empirical results have demonstrated the efficacy of our proposed debiased method in mitigating source bias to different extents by adjusting the debiased coefficient $\alpha$.
+This flexibility allows for customizing debiasing mechanisms to meet diverse perspectives and demands.
 Notably, the decision to maintain equality between the two content sources or favor human-written content can be tailored based on specific requirements and environmental considerations.
-The optimal strategy for enhancing the IR ecosystem remains an open question that warrants further exploration.
+For example, users may not mind the content’s source if it is of high quality and fulfills their informational needs.
+However, bias becomes a significant issue when we aim to credit content providers and encourage more creation, impacting the sustainability of the content creation ecosystem.
+The optimal strategy for enhancing the sustainable development of the IR ecosystem remains an open question for further exploration.
 6. Discussion: Sounding the Alarm
-In this paper, we have sought to extrapolate the performance of various retrieval models in the LLM era, where the corpus encompasses both human-written and LLM-generated content.
 Through a rigorous series of experiments and thorough analysis, we have identified that neural retrieval models demonstrate clear preferences for LLM-generated texts, referred to as source bias.
 This bias, with the burgeoning proliferation of LLMs and AIGC, may raise significant concerns for a variety of aspects.
-First, the source bias has the potential to disrupt the ecosystem of information retrieval.
-On the one hand, our findings underscore the growing dominance of LLM-generated text in IR systems, which makes human-written content gradually unsearchable and could diminish the incentive for content providers to continue creating human-written content.
-On the other hand, users may find themselves dissatisfied with search results that predominantly consist of LLM-generated content (Cao et al., 2023; Wu et al., 2023a; Chen et al., 2023), leading to a high user churn rate of the IR systems.
-Consequently, the benefits and experiences of all stakeholders within the IR ecosystem, including platforms, content providers, and users, may erode over time.
-Second, the source bias might inadvertently amplify the spread of LLM-generated misinformation.
-Some recent studies have underscored the ability of LLMs (e.g., ChatGPT) to generate misleading or deceptive information, either unintentionally or intentionally (Su et al., 2023; Chen and Shu, 2023; Pan et al., 2023).
-Notably, LLM-generated misinformation is proliferating at an alarming pace, outpacing human-generated misinformation, particularly on mainstream news websites (Hanley and Durumeric, 2023; Spitale et al., 2023).
-What’s more, our findings directly suggest that neural retrieval models tend to favor LLM-generated text, which means that LLM-generated misinformation may occupy higher positions in ranking lists in current IR systems.
-As a result, the source bias can exacerbate the dissemination of LLM-generated misinformation, potentially posing significant social issues.
-Finally, the source bias could be exploited intentionally as a means to manipulate existing retrieval and ranking algorithms.
-Malicious actors are more likely to leverage LLMs to generate text with a higher likelihood of ranking at a more advantageous position.
-Indeed, our empirical investigations reveal that malicious users can easily instruct LLMs to generate text with similar semantic information, e.g., through prompts for paraphrase generation and rewriting generation.
-Thus, this bias could be weaponized as a potential method of attack against neural retrieval models within today’s search engines, reminiscent of earlier web spam link attacks against PageRank  (Gyöngyi et al., 2004).
+First, the presence of source bias poses a significant risk of gradually rendering human-written content less accessible, potentially causing a disruption in the content ecosystem.
+More severely, the concern is escalating with the growing prevalence of LLM-generated content online(Hanley and Durumeric, 2023; Bengio et al., 2023).
+Second, there is the risk that source bias may amplify the spread of misinformation, especially considering the potential of LLMs to generate deceptive content, whether intentionally or not (Chen and Shu, 2023; Su et al., 2023; Pan et al., 2023; Aslett et al., 2023).
+Third, source bias may be maliciously exploited to attack against neural retrieval models within today’s search engines, creating a precarious vulnerability that could be weaponized by malicious actors, reminiscent of earlier web spam link attacks against PageRank (Gyöngyi et al., 2004).
 As discussed above, since LLMs can be readily instructed to generate texts at scale, source bias presents potential tangible and serious threats to the ecosystem of web content, public trust, and online safety.
-We hope this discussion will raise public awareness and sound the alarm regarding the risks posed by source bias in the LLM era.
+We hope this discussion will sound the alarm regarding the risks posed by source bias in the LLM era.
 7. Related Work
-7.1. Large Language Models for IR
-The emergence of large language models (LLMs) (Zhao et al., 2023) has ushered in a transformative era across various research domains, such as natural language processing (NLP) (Bubeck et al., 2023; Bang et al., 2023), education (Guo et al., 2023; Nunes et al., 2023), recommender systems (Dai et al., 2023; Fan et al., 2023), finance (Wu et al., 2023b; Huang et al., 2023), and medicine (Alberts et al., 2023; Thirunavukarasu et al., 2023).
+Large Language Models for IR.
+The emergence of large language models (LLMs) (Zhao et al., 2023b; Wei et al., 2022; Yang et al., 2023) has ushered in a transformative era across various research domains, such as natural language processing (NLP) (Bubeck et al., 2023; Bang et al., 2023), education (Guo et al., 2023; Nunes et al., 2023), recommender systems (Dai et al., 2023; Fan et al., 2023), finance (Wu et al., 2023b; Huang et al., 2023), and medicine (Alberts et al., 2023; Thirunavukarasu et al., 2023).
 In the field of IR, much effort has also been made to utilize the remarkable knowledge and capabilities of LLMs to enhance IR systems (Ai et al., 2023; Zhu et al., 2023).
-In the industry community, an exemplary successful application is New Bing [^5] https://www.bing.com/new, which is an LLM-powered search assistant that adeptly extracts information from various web pages and delivers concise summarized responses to user queries, thereby elevating the overall search experience.
-In the research community, there has been a proactive exploration of integrating LLMs into the major IR components, including query rewriters (Srinivasan et al., 2022; Wang et al., 2023a), retrievers (Xu et al., 2022; Dai et al., 2022), re-rankers (Sun et al., 2023; Cho et al., 2023), and readers (Shi et al., 2023; Izacard et al., 2022).
-For a more comprehensive overview of the recent advancements in the LLMs for IR, we recommend referring to the recent survey (Zhu et al., 2023).
-7.2. Artificial Intelligence Generated Content
-Artificial Intelligence Generated Content (AIGC) is a rapidly advancing field that involves the creation of content using advanced Generative AI (GAI) techniques (Cao et al., 2023; Wu et al., 2023a; Agnese et al., 2020).
+In the industry community, an exemplary successful application is New Bing [^6] https://www.bing.com/new, which is an LLM-powered search assistant that adeptly extracts information from various web pages and delivers concise summarized responses to user queries, thereby improving the search experience.
+In the research community, there has been a proactive exploration of integrating LLMs into the IR components, including query rewriters (Srinivasan et al., 2022; Wang et al., 2023a), retrievers (Xu et al., 2022; Dai et al., 2022), re-rankers (Sun et al., 2023; Cho et al., 2023), and readers (Shi et al., 2023; Izacard et al., 2022).
+For a more comprehensive overview of the recent advancements in LLMs for IR, please refer to the recent survey (Zhu et al., 2023).
+Artificial Intelligence Generated Content.
+Artificial Intelligence Generated Content (AIGC) is a rapidly advancing field that involves the creation of content using advanced Generative AI (GAI) (Cao et al., 2023; Wu et al., 2023a; Agnese et al., 2020).
 Unlike traditional content crafted by humans, AIGC can be generated at scale and in considerably less time (Hanley and Durumeric, 2023; Spitale et al., 2023).
 Recently, the development of LLMs and other GAI models has greatly improved the quality of AIGC content than before.
-For instance, LLMs such as ChatGPT have shown impressive emergent abilities in generating human-like content (Cao et al., 2023; Wu et al., 2023a).
+For instance, LLMs such as ChatGPT have shown impressive abilities in generating human-like content (Cao et al., 2023; Wu et al., 2023a).
 The DALL-E-3 (Betker et al., 2023), another state-of-the-art text-to-image generation system, can follow user instructions to produce high-quality images.
-Nevertheless, as AIGC becomes more prevalent across myriad domains, ethical concerns and potential risks come into sharper focus (Thakur et al., 2023; Wang et al., 2023b).
+Nevertheless, as AIGC becomes more prevalent across myriad domains, ethical concerns, and potential risks come into sharper focus (Thakur et al., 2023; Wang et al., 2023b).
 In fact, inevitably, the GAI models may generate content with bias and discrimination as the large training data always contain bias and toxicity (Zhuo et al., 2023; Deshpande et al., 2023; Bengio et al., 2023).
@@
 Finally, we discuss the crucial concerns and potential risks of this bias to the whole web ecosystem.
-Our findings provide valuable insights into several promising directions for future research.
-Firstly, under siege by the LLMs, how to mitigate the observed source bias and ensuring the sustainable development of information ecosystem is a crucial research direction.
-Secondly, although our focus in this paper centers on information retrieval (IR) systems, the findings and conclusion are not IR-specific.
-An interesting research question to explore further is whether this bias also manifests in other information systems, such as recommender and advertising systems.
-Thirdly, it’s essential to recognize that neural models may exhibit biases not only towards generated text but also other data modalities, such as images and multimodal content.
-Exploring the source bias on tasks beyond texts, such as image-text retrieval, especially in light of emerging generative models like DALL-E-3 (Betker et al., 2023) capable of generating high-quality images, represents a promising research direction.
-Appendix A Appendix
-A.1. Example of Evaluation Metrics
+Our study offers valuable insights into several promising directions for future research, including exploring source bias in other information systems (e.g., recommender systems and advertising systems) and examining source bias in neural models towards AIGC data across multiple data modalities, not limited to text.
+Moreover, uncovering the root cause of the source bias and thus further mitigating it are difficult but crucial research directions.
+Acknowledgements.
+This work was funded by the National Key R&D Program of China (2023YFA1008704), the National Natural Science Foundation of China (No.
+62377044, 62276248, 62376275, 62076234), Beijing Natural Science Foundation (No.
+4222029), Beijing Key Laboratory of Big Data Management and Analysis Methods, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, PCC@RUC, funds for building world-class universities (disciplines) of Renmin University of China, and the Youth Innovation Promotion Association CAS under Grants No.2023111.
+This work was supported by the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China (RUC24QSDL013).
+We thank all the anonymous reviewers for their positive and insightful comments.
+Appendix A Example of Evaluation Metrics
+Figure 10.
+A toy example to illustrate how ranking metrics are calculated for each target corpus.
+In this toy example, given a query, the top 6 documents are retrieved and the rank is from left to right in descending order.
+Two relevant documents (i.e., $d_{1}^{G}$ and $d_{1}^{H}$) are highlighted with red boxes.
 In this section, we provide a toy example for illustrating the calculation of the evaluation metrics for exploring source bias, as depicted in Figure 10.
-Note that when calculating metrics for one target corpus, labels of the document from another are masked as negative.
+Specifically, for each query, an IR model produces a ranking list that comprises documents from both human-written and LLM-generated corpora.
+We then calculate ranking metrics separately for human-written and LLM-generated texts, depending on the target data source.
+When computing metrics for one target (e.g., human-written corpus $\mathcal{C}^{H}$), the data corresponding to the other side (e.g., LLM-generated corpus $\mathcal{C}^{G}$) is disregarded (i.e., mask all the positive label $r\in\mathcal{R}^{G}$ as negative), but the rank of each document is still based on the original ranking list that incorporates a mixture of both types of text.
 For instance, in this toy example, when targeting the human-written corpus, the relevant document $d_{1}^{G}$ generated by LLM is treated as a negative sample.
-And when calculating the ranking metrics, only consider the rank of the positive human-written documents.
+And when calculating the ranking metrics, we only consider the rank of the positive human-written documents.
 When the target corpus is LLM-generated, we adopt the same principle, i.e., only take the rank of the positive LLM-generated documents into account for calculating the ranking metrics.
-A.2. Implementation Details
-We call the OpenAI API gpt-3.5-turbo-0613 for ChatGPT and llama-2-7b-chat model for LLaMA2  [^6] https://github.com/facebookresearch/llama.
-If not specified, the temperature of the used LLM for generation corpus is set as 0.2 by default for the experiments.
-In all our experiments, the nucleus sampling ratio (i.e., top_p) is 1 for both ChatGPT and Llama2.
-For the LLM-generated corpus, we conduct post-processing to remove unrelated parts of the original response from LLM like “Sure, here’s a possible rewrite of the text:”.
-We follow the standard practice in BEIR (Thakur et al., 2021) benchmark to evaluate all the retrieval models.
-All the models are trained on MS MARCO (Nguyen et al., 2016).
-For the re-ranking models, we re-rank the top-$100$ documents retrieved by BM25.
-A.3. Proof of Theorem 4.1
-In this section, we provided the detailed proof of Theorem 4.1
-Proof.
+Appendix B More Experimental Results
+In Table 7, we provide the results for source bias with more common prompts sourced from InstructGPT-prompts Github Repository  [^7] https://github.com/kevinamiri/Instructgpt-prompts#rephrase-a-passage.
+These results indicate that common prompts can easily trigger source bias with LLM-generated content.
+Table 7.
+Overall source bias in neural retrievers w.r.t.
+Relative $\Delta$ (NDCG@1) on SciFact+AIGC with mixed human-written and Llama2-generated corpora generated from different common rephrasing prompts.
+Prompt ANCE BERM TAS-B Contriever
+Rewrite the text below in your own words: -7.0 -22.8 -11.2 -27.6
+Paraphrase the provided text while maintaining its meaning: -26.0 -55.3 -24.1 -13.6
+Summarize the following passage in a concise manner: -1.4 -43.3 -34.0 -32.4
+Simplify the given passage while keeping the main ideas: -29.0 -21.7 -22.5 -40.9
+Rephrase the given text using alternative expressions: -25.3 -34.7 -61.9 -18.4
+Condense the following passage to focus on key points: -19.0 -24.8 -22.8 -16.4
+Briefly restate the provided text without losing its essence: -29.6 -50.7 -41.3 -29.8
+Reword the passage below to make it more succinct: -40.6 -71.1 -54.7 -34.0
+Express the following text in a different way while keeping its intent: -50.7 -39.8 -34.9 0.0
+Appendix C Theoretical Analysis and Insights
+In  Figure 7, we have compared the PPL for different corpus using the BERT model.
+In this section, we aim to further provide some theoretical insights into the above observations that LLM-generated texts have a smaller perplexity than human-written texts.
+Without loss of generality, we define the PPL in an autoregressive manner.
+Let $d^{H}$ denote a document written by humans, and $d^{G}$ a document generated by an LLM conditioned on $d^{H}.$ For a given document $d$ and BERT model $\mathcal{B}$, PPL is calculated as
+$\text{PPL}(d,\mathcal{B})=-\frac{1}{S}\sum_{s=1}^{S}\log P_{\text{BERT}}(d_{s}|d_{<{s}}).$
+Similarly, we use $\text{PPL}(d,\mathcal{H})$ to represent the PPL of document $d$ when evaluated by humans. The PPL of $d^{G}$ conditioned on $d^{H}$ is denoted as
+$\text{PPL}(d^{G}\mid d^{H},\mathcal{B})=-\frac{1}{S}\sum_{s=1}^{S}\log P_{\text{BERT}}(d_{s}^{G}|d_{<{s}}^{G},d^{H}).$
+When evaluated by humans, we use $\text{PPL}(d^{G}\mid d^{H},\mathcal{H})$ to represent the PPL of $d^{G}$ conditioned on $d^{H}$.
+In the theorem below, we introduce three assumptions: Semantic Superiority, Conditional Redundancy, and Bounded Perplexity, to theoretically establish the sufficient conditions under which $\text{PPL}(d^{G},\mathcal{B})\leq\text{PPL}(d^{H},\mathcal{B})$ holds.
+Semantic Superiority suggests that the perplexity of human-written texts, when evaluated by humans, is lower than when evaluated by BERT.
+Conditional Redundancy implies that the perplexity of $d^{G}$, given $d^{H}$, is lower than the perplexity of $d^{H}$ when evaluated directly.
+This is intuitively true when the information added in generating from $d^{H}$ to $d^{G}$ doesn’t exceed the original information in $d^{H}$.
+Bounded perplexity assumes that there exists an upper bound $\epsilon$ on the increase in perplexity when evaluating $d^{G}$directly, compared to evaluating $d^{G}$ conditioned on $d^{H}$.
+Then we have the following theorem:
+Theorem C.1.
 Given the following conditions:
-  • [leftmargin=0.5cm]
-  • Semantic Superiority: human beings outperform BERT in understanding human-authored texts, i.e.,
+  • Semantic Superiority: human beings outperform BERT in understanding human-written texts, i.e.,
     $\textnormal{PPL}(d^{H},\mathcal{B})-\textnormal{PPL}(d^{H},\mathcal{H})\geq 0.$
@@
     $\textnormal{PPL}(d^{H},\mathcal{H})-\textnormal{PPL}(d^{G}\mid d^{H},\mathcal{H})\geq 0.$
-  • Bounded Perplexity: there exists non-negative value $\epsilon$ such that
+  • Bounded Perplexity: there exists a bounded non-negative difference $\epsilon$ in BERT’s perplexity for $d^{G}$ with or without $d^{H},$ i.e.,
     $\textnormal{PPL}(d^{G},\mathcal{B})-\textnormal{PPL}(d^{G}\mid d^{H},\mathcal{B})\leq\epsilon.$
-Our goal is to prove $\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\left[\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{H},\mathcal{B})\right]\leq 0$ under the condition that LLMs are closer to BERT than to humans when predicting $d^{G}$ given $d^{H}.$ This condition states that, for any $s\in[S],$ the KL divergences satisfy
-$D_{\textnormal{KL}}(P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})\|P_{\text{BERT}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})+\epsilon$
-$\leq D_{\textnormal{KL}}(P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})\|P_{\text{Human}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})).$
+If LLM aligns more closely with BERT than with humans when predicting $d^{G}$ given $d^{H}$, such that for any $s\in[S],$
+(3) $D_{\textnormal{KL}}\left(P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})\|P_{\text{BERT}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})\right)+\epsilon$
+$\leq D_{\textnormal{KL}}\left(P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})\|P_{\text{Human}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})\right),$
+it follows that
+$\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\left[\textnormal{PPL}(d^{G},\mathcal{B})-\textnormal{PPL}(d^{H},\mathcal{B})\right]\leq 0.$
+In  Theorem C.1, the KL divergence is used to compare the distributions of the document $d^{G}$ conditioned on $d^{H}$ according to the LLM, BERT model, and humans.
+It is worth emphasizing that inequation  (LABEL:eqn:KL-distance) is not the assumption on the understanding capabilities of BERT, LLM, and humans.
+Instead, this inequation assumes that when predicting $d^{G}$ given $d^{H},$ the predictions by LLM are more closely aligned with those of BERT.
+We demonstrate that, when inequation  (LABEL:eqn:KL-distance) is satisfied, the perplexity (evaluated by PLMs such as BERT) of $d^{G}$ is lower than that of $d^{H}$.
+We’d like to emphasize that it is reasonable to expect that inequation  (LABEL:eqn:KL-distance) holds true because both LLM and BERT are Transformer-based models that use similar pretraining paradigms.
+The commonality in model structure and learning paradigms may lead to similar inherent biases in text prediction, making their predictions more aligned with each other.
+The proof for  Theorem C.1 is provided as follows:
+Proof.
 We start by introducing the term $\text{PPL}(d^{H},\mathcal{H}):$
@@
 PPL $(d^{G},\mathcal{B})-\text{PPL}(d^{H},\mathcal{B})\leq\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{H},\mathcal{H})$
-$=\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{L})+\text{PPL}(d^{G}\mid d^{H},\mathcal{L})-\text{PPL}(d^{H},\mathcal{H}).$
-Next, we provide upper bounds of $\text{PPL}(d^{G}\mid d^{H},\mathcal{L})-\text{PPL}(d^{H},\mathcal{H}):$
-$\text{PPL}(d^{G}\mid d^{H},\mathcal{L})-\text{PPL}(d^{H},\mathcal{H})$
-$=\text{PPL}(d^{G}\mid d^{H},\mathcal{L})-\text{PPL}(d^{G}\mid d^{H},\mathcal{H})+\text{PPL}(d^{G}\mid d^{H},\mathcal{H})-\text{PPL}(d^{H},\mathcal{H})$
-$\leq\text{PPL}(d^{G}\mid d^{H},\mathcal{L})-\text{PPL}(d^{G}\mid d^{H},\mathcal{H}),$
-where the inequality follows from the Conditional Redundancy. Taking expectation on $\text{PPL}(d^{G}\mid d^{H},\mathcal{L})-\text{PPL}(d^{G}\mid d^{H},\mathcal{H}):$
-$S\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\left[\text{PPL}(d^{G}\mid d^{H},\mathcal{L})-\text{PPL}(d^{G}\mid d^{H},\mathcal{H})\right]$
-$=-\sum_{s=1}^{S}\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\log\frac{P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{L},d^{H})}{P_{\text{Human}}(d_{s}^{G}|d_{<{s}}^{L},d^{H})}$
-$=-\sum_{s=1}^{S}\mathbb{E}_{P_{\text{LLM}}(d_{<s}^{G}\mid d^{H})}\mathbb{E}_{P_{\text{LLM}}(d_{s}^{G}\mid d_{<s}^{G},d^{H})}\log\frac{P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{L},d^{H})}{P_{\text{Human}}(d_{s}^{G}|d_{<{s}}^{L},d^{H})}$
-$=-\sum_{s=1}^{S}\mathbb{E}_{P_{\text{LLM}}(d_{<s}^{G}\mid d^{H})}\underbrace{D_{\textnormal{KL}}(P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{L},d^{H})\|P_{\text{Human}}(d_{s}^{G}|d_{<{s}}^{L},d^{H}))}_{D_{\textnormal{KL}}(P_{\text{LLM}}\|P_{\text{Human}})}.$
-Similarly, $\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{L})$ can be rewritten as:
-$\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{L})$
-$=\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{B})+\text{PPL}(d^{G}\mid d^{H},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{L}).$
-Taking expectation on $\text{PPL}(d^{G}\mid d^{H},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{L}):$
-$S\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\left[\text{PPL}(d^{G}\mid d^{H},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{L})\right]$
-$=\sum_{s=1}^{S}\mathbb{E}_{P_{\text{LLM}}(d_{<s}^{L}\mid d^{H})}\underbrace{D_{\textnormal{KL}}(P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{L},d^{H})\|P_{\text{BERT}}(d_{s}^{G}|d_{<{s}}^{L},d^{H}))}_{D_{\textnormal{KL}}(P_{\text{LLM}}\|P_{\text{BERT}})}.$
+$=\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{G})$
+$\quad+\text{PPL}(d^{G}\mid d^{H},\mathcal{G})-\text{PPL}(d^{H},\mathcal{H}).$
+Next, we provide upper bounds of $\text{PPL}(d^{G}\mid d^{H},\mathcal{G})-\text{PPL}(d^{H},\mathcal{H}):$
+$\text{PPL}(d^{G}\mid d^{H},\mathcal{G})-\text{PPL}(d^{H},\mathcal{H})$
+$=\text{PPL}(d^{G}\mid d^{H},\mathcal{G})-\text{PPL}(d^{G}\mid d^{H},\mathcal{H})$
+$\quad+\text{PPL}(d^{G}\mid d^{H},\mathcal{H})-\text{PPL}(d^{H},\mathcal{H})$
+$\leq\text{PPL}(d^{G}\mid d^{H},\mathcal{G})-\text{PPL}(d^{G}\mid d^{H},\mathcal{H}),$
+where the inequality follows from the Conditional Redundancy. Taking expectation on $\text{PPL}(d^{G}\mid d^{H},\mathcal{G})-\text{PPL}(d^{G}\mid d^{H},\mathcal{H}):$
+$-S\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\left[\text{PPL}(d^{G}\mid d^{H},\mathcal{G})-\text{PPL}(d^{G}\mid d^{H},\mathcal{H})\right]$
+$=\sum_{s=1}^{S}\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\log\frac{P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})}{P_{\text{Human}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})}$
+$=\sum_{s=1}^{S}\mathbb{E}_{P_{\text{LLM}}(d_{<s}^{G}\mid d^{H})}\mathbb{E}_{P_{\text{LLM}}(d_{s}^{G}\mid d_{<s}^{G},d^{H})}\log\frac{P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})}{P_{\text{Human}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})}$
+$=\sum_{s=1}^{S}\mathbb{E}_{P_{\text{LLM}}(d_{<s}^{G}\mid d^{H})}\underbrace{D_{\textnormal{KL}}(P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})\|P_{\text{Human}}(d_{s}^{G}|d_{<{s}}^{G},d^{H}))}_{D_{\textnormal{KL}}(P_{\text{LLM}}\|P_{\text{Human}})}.$
+Similarly, $\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{G})$ can be rewritten as:
+$\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{G})$
+$=\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{B})$
+$\quad+\text{PPL}(d^{G}\mid d^{H},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{G}).$
+Taking expectation on $\text{PPL}(d^{G}\mid d^{H},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{G}):$
+$S\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\left[\text{PPL}(d^{G}\mid d^{H},\mathcal{B})-\text{PPL}(d^{G}\mid d^{H},\mathcal{G})\right]$
+$=\sum_{s=1}^{S}\mathbb{E}_{P_{\text{LLM}}(d_{<s}^{L}\mid d^{H})}\underbrace{D_{\textnormal{KL}}(P_{\text{LLM}}(d_{s}^{G}|d_{<{s}}^{G},d^{H})\|P_{\text{BERT}}(d_{s}^{G}|d_{<{s}}^{G},d^{H}))}_{D_{\textnormal{KL}}(P_{\text{LLM}}\|P_{\text{BERT}})}.$
 Thus,
@@
 $\mathbb{E}_{P_{\text{LLM}}(d^{G}\mid d^{H})}\left[\text{PPL}(d^{G},\mathcal{B})-\text{PPL}(d^{H},\mathcal{B})\right]\leq\epsilon-\epsilon=0.$
-Figure 12.
-A toy example to illustrate how ranking metrics are calculated for each target corpus.
-In this toy example, given a query, the top 6 documents are retrieved and the rank is from left to right in descending order.
-Two relevant documents (i.e., $d_{1}^{G}$ and $d_{1}^{H}$) are highlighted with red boxes.
-(a) Human vs. ChatGPT (T=0.2)
-(b) Human vs. Llama2 (T=0.2)
-Figure 13. Semantic embedding visualization of human-written and LLM-generated corpora on SciFact+AIGC datasets.
-A.4. More Results of Dataset Quality Validation
-In Section 3.2 and 3.5, we also leverage ChatGPT (Temperature=0.2) and Llama2 (Temperature=1.0) to generate LLM-generated corpus for further experiments.
-For data quality evaluation, we also analyze the semantic similarity.
-Figure 13 and Figure 14 show the visualization of T-SNE and cosine similarity of the sementic embedding from text-embedding-ada-002777https://platform.openai.com/docs/api-reference/embeddings, respectively.
-As we can see, the LLM-generated texts almost have the same semantic information of the corresponding original human-written texts.
-We also conduct experiments with different retrieval models on the sole corpus.
-As depicted in Table 7 and Table 8, there are no significant performance discrepancies in terms of various ranking metrics between the human-written and LLM-generated corpora, further affirming the quality of our newly constructed datasets.
-A.5. Raw Results with Llama2 (T=1.0)
-The raw results of retrieval models for the mixed human-written and Llama2-generated corpora ($T=1.0$) are shown in Table 9.
-Once again, the results highlight the presence of source bias in both neural retrievers and re-rankers.
-A.6. Examples of LLM-Generated Corpus
-We provide examples of human-written corpus and the corresponding rewritten corpus by Llama2 and ChatGPT on SciFact+AIGC and NQ320K+AIGC dataset, as shown in  subsection A.6 and  subsection A.6, respectively.
-For different examples in each dataset, we indicate the LLM model and temperature in the sampling strategy.
-(a) Llama2 (T=0.2) vs. ChatGPT (T=0.2)
-(b) Llama2 (T=0.2) vs. Llama2 (T=1.0)
-Figure 14. Distribution of cosine similarity of semantic embedding between LLM-generated and human-written corpora.
-Table 7. Performance comparison of different retrieval models on the sole human-written corpus or ChatGPT-generated corpus on SciFact+AIGC dataset. The temperature of ChatGPT is set as 0.2.
-Model Type Model Corpus NDCG@1 NDCG@3 NDCG@5 MAP@1 MAP@3 MAP@5
-Lexical TF-IDF Human-Written 42.0 49.5 52.7 40.7 47.1 49.0
-LLM-Generated 41.7 49.9 52.1 39.4 47.1 48.6
-BM25 Human-Written 46.0 54.2 56.3 43.8 51.5 52.8
-LLM-Generated 46.7 53.9 56.0 44.4 51.5 52.7
-Neural ANCE Human-Written 38.7 44.3 46.5 36.3 41.9 43.3
-LLM-Generated 40.3 45.5 47.9 38.2 43.3 44.9
-BERM Human-Written 37.0 42.1 44.2 34.7 39.7 41.3
-LLM-Generated 37.3 43.2 45.0 35.1 40.8 42.0
-TAS-B Human-Written 52.7 58.1 60.2 49.9 55.6 57.2
-LLM-Generated 50.0 57.3 59.8 48.2 54.9 56.5
-Contriever Human-Written 54.0 61.8 63.2 51.4 58.9 60.0
-LLM-Generated 55.7 62.1 64.8 53.3 59.7 61.5
-Table 8. Performance comparison of different retrieval models on the sole human-written corpus or Llama2-generated corpus on SciFact+AIGC datasets. The temperature of Llama2 is set as 1.0.
-Model Type Model Corpus NDCG@1 NDCG@3 NDCG@5 MAP@1 MAP@3 MAP@5
-Lexical TF-IDF Human-Written 42.0 49.5 52.7 40.7 47.1 49.0
-LLM-Generated 42.0 48.9 50.7 39.8 46.4 47.5
-BM25 Human-Written 46.0 54.2 56.3 43.8 51.5 52.8
-LLM-Generated 46.0 52.4 54.7 43.6 50.0 51.5
-Neural ANCE Human-Written 38.7 44.3 46.5 36.3 41.9 43.3
-LLM-Generated 39.7 44.9 46.5 36.9 42.6 43.7
-BERM Human-Written 37.0 42.1 44.2 34.7 39.7 41.3
-LLM-Generated 38.0 42.3 43.5 35.4 40.0 41.0
-TAS-B Human-Written 52.7 58.1 60.2 49.9 55.6 57.2
-LLM-Generated 52.3 55.5 59.0 49.5 53.7 55.9
-Contriever Human-Written 54.0 61.8 63.2 51.4 58.9 60.0
-LLM-Generated 52.0 59.0 62.0 49.6 56.4 58.4
-Table 9. Performance comparison of different retrieval models for mixed human-written and Llama2-generated corpora on SciFact+AIGC dataset. The temperature of Llama2 is set as 1.0.
-Model Type Model Corpus NDCG@1 NDCG@3 NDCG@5 MAP@1 MAP@3 MAP@5
-Lexical TF-IDF Human-Written 21.7 36.5 39.3 20.9 32.6 34.3
-LLM-Generated 15.3 32.8 35.1 14.5 28.2 29.6
-$\text{Relative}~{}~{}\Delta$ 34.6 10.7 11.3 36.2 14.5 14.7
-BM25 Human-Written 24.3 39.7 43.2 23.2 35.6 37.6
-LLM-Generated 22.7 39.4 41.6 21.4 35.0 36.3
-$\text{Relative}~{}~{}\Delta$ 6.8 0.8 3.8 8.1 1.7 3.5
-Neural ANCE Human-Written 15.7 30.8 33.1 14.1 26.7 28.1
-LLM-Generated 23.7 34.2 36.6 22.6 31.2 32.6
-$\text{Relative}~{}~{}\Delta$40.6 -10.5 -10.0 -46.3 -15.5 -14.8
-BERM Human-Written 16.7 29.6 31.8 15.9 26.3 27.6
-LLM-Generated 22.3 32.1 33.9 20.4 29.0 30.1
-$\text{Relative}~{}~{}\Delta$28.7 -8.1 -6.4 -24.8 -9.8 -8.7
-TAS-B Human-Written 21.0 41.0 43.2 20.0 35.9 37.2
-LLM-Generated 32.3 44.5 46.8 30.7 41.1 42.4
-$\text{Relative}~{}~{}\Delta$42.4 -8.2 -8.0 -42.2 -13.5 -13.1
-Contriever Human-Written 25.0 43.9 47.8 24.5 39.3 41.6
-LLM-Generated 29.0 44.4 47.4 27.4 40.2 41.9
-$\text{Relative}~{}~{}\Delta$14.8 -1.1 0.8 -11.2 -2.3 -0.7
-Re-ranking BM25+MiniLM Human-Written 22.0 42.0 47.0 21.5 37.3 39.8
-LLM-Generated 31.0 47.0 50.0 29.3 42.2 43.9
-$\text{Relative}~{}~{}\Delta$34.0 -11.2 -6.2 -30.7 -12.3 -9.8
-BM25+monoT5 Human-Written 24.7 47.1 50.5 23.8 41.5 43.6
-LLM-Generated 35.0 51.6 53.8 32.9 47.0 48.4
-$\text{Relative}~{}~{}\Delta$34.5 -9.1 -6.3 -32.1 -12.4 -10.4
-Table 10. Examples of human-written and LLM-generated texts on SciFact+AIGC dataset.
-Human
-Likelihood ratios are one of the best measures of diagnostic accuracy, although they are seldom used, because interpreting them requires a calculator to convert back and forth between “probability” and “odds” of disease.
-This article describes a simpler method of interpreting likelihood ratios, one that avoids calculators, nomograms, and conversions to “odds” of disease.
-Several examples illustrate how the clinician can use this method to refine diagnostic decisions at the bedside.
-[…]
-Llama2 (Temperature=0.2)
-Likelihood ratios are a valuable tool for assessing the accuracy of diagnostic tests, yet they are often overlooked due to the complexity of interpreting them.
-Converting between "probability" and "odds" of disease requires a calculator, which can be a barrier for clinicians at the bedside.
-This article proposes a simpler method of interpreting likelihood ratios, eliminating the need for calculators, nomograms, or conversions to "odds" of disease.
-With this approach, clinicians can make more informed diagnostic decisions in real-time.
-Several examples demonstrate how this method can be applied in practice.
-[…]
-Llama2 (Temperature=1.0)
-Likelihood ratios are a powerful tool for evaluating the accuracy of diagnostic tests, yet they are often overlooked due to the complexity of interpreting them.
-Conversion between "probability" and "odds" of disease requires a calculator, which can make the process challenging for clinicians.
-This article presents a simplified approach to interpreting likelihood ratios, eliminating the need for calculators, nomograms, or conversions to "odds" of disease.
-By following this method, clinicians can improve their diagnostic decision-making at the bedside through practical examples.
-[…]
-ChatGPT (Temperature=0.2)
-The use of likelihood ratios is considered one of the most effective ways to assess diagnostic accuracy.
-However, they are not commonly utilized due to the complexity of interpreting them, which involves converting between the "probability" and "odds" of disease using a calculator.
-This article presents a simpler approach to interpreting likelihood ratios that eliminates the need for calculators, nomograms, and conversions to disease odds.
-Through various examples, it demonstrates how clinicians can employ this method to enhance diagnostic decision-making directly at the patientś bedside.
-Table 11. Examples of human-written and LLM-generated texts on NQ320K+AIGC dataset.
-Case 1
-[…]
-2018 FIFA World Cup disciplinary record In the 2018 FIFA World Cup , which took place in Russia from 14 June to 15 July 2018 , the main disciplinary actions taken against players came in the form of red and yellow cards .
-A player sanctioned with a red card was sent off from the pitch and could not be replaced .
-Furthermore , the player was automatically banned from his team ’s next match .
-After a straight red card , FIFA conducted a hearing and considered extending this ban beyond one match .
-If the ban extended beyond the end of the World Cup finals ( for example , a player was sent off in his team ’s last match ) , it had to be served in the team ’s next competitive international match ( es ) .
-A player also received a one match ban if he picked up a single yellow card in two different games from the beginning of the group stage until the end of the quarter - finals , where single yellow cards were negated .
-However , such a ban did not carry over beyond the World Cup finals if the
-[…]
-Llama2 (Temperature=0.2)
-In the 2018 FIFA World Cup, which took place in Russia from June 14 to July 15, 2018, the main disciplinary actions against players were red and yellow cards.
-A player sent off with a red card was unable to be replaced, and was automatically banned from their team’s next match.
-Following a straight red card, FIFA conducted a hearing and could have extended this ban beyond one match.
-If the ban extended beyond the end of the World Cup finals (e.g., a player was sent off in their team’s last match), it had to be served in their team’s next competitive international match.
-A player received a one-match ban if they accumulated a single yellow card in two or more games from the start of the group stage to the end of the quarter-finals, where single yellow cards were voided.
-However, this ban did not carry over beyond the World Cup finals if the player’s team was eliminated before the quarter-finals.
-Case 2
-[…]
-San Andreas fault The San Andreas Fault is a continental transform fault that extends roughly 1,200 kilometers ( 750 mi ) through California .
-It forms the tectonic boundary between the Pacific Plate and the North American Plate , and its motion is right - lateral strike - slip ( horizontal ) .
-The fault divides into three segments , each with different characteristics and a different degree of earthquake risk , the most significant being the southern segment , which passes within about 35 miles ( 56 km ) of Los Angeles .
-The slip rate along the fault ranges from 20 to 35 mm ( 0.79 to 1.38 in ) / yr .
-The fault was first identified in 1895 by Professor Andrew Lawson of UC Berkeley , who discovered the northern zone .
-It is often described as having been named after San Andreas Lake , a small body of water that was formed in a valley between the two plates .
-However , according to some of his reports from 1895 and 1908 , Lawson actually named it after the surrounding San Andreas Valley .
-Following the 1906 San Francisco earthquake , Lawson concluded that the
-[…]
-Llama2 (Temperature=0.2)
-The San Andreas Fault is a continental transform fault that spans approximately 1,200 kilometers (750 miles) through California, serving as the tectonic boundary between the Pacific Plate and the North American Plate.
-The fault’s motion is right-lateral strike-slip (horizontal), causing the plates to move horizontally past each other.
-The fault is divided into three segments, each with distinct characteristics and varying degrees of earthquake risk, with the southern segment posing the greatest risk, passing within approximately 35 miles (56 km) of Los Angeles.
-The slip rate along the fault ranges from 20 to 35 mm (0.79 to 1.38 in) per year.
-The fault was first identified in 1895 by Professor Andrew Lawson of UC Berkeley, who discovered the northern zone.
-Contrary to popular belief, the fault was not named after San Andreas Lake, a small body of water located in a valley between the two plates.
-Instead, Lawson actually named it after the surrounding San Andreas Valley, as revealed in his reports from 1895 and 1908.
-Following the 1906 San Francisco earthquake, Lawson concluded that the fault was responsible for the devastating event.